### v1

# On the stability of global forecasting models

Abstract
Forecast stability, that is the consistency of predictions over time, is essential in business settings where sudden shifts in forecasts can disrupt planning and erode trust in predictive systems. Despite its importance, stability is often overlooked in favor of accuracy, particularly in global forecasting models. In this study, we evaluate the stability of point and probabilistic forecasts across different retraining frequencies and ensemble strategies using two large retail datasets (M5 and VN1). To do this, we introduce a new metric for probabilistic stability (MQC) and analyze ten different global models and four ensemble configurations. The results show that less frequent retraining not only preserves but often improves forecast stability, while ensembles, especially those combining diverse pool of models, further enhance consistency without sacrificing accuracy. These findings challenge the need for continuous retraining and highlight ensemble diversity as a key factor in reducing forecast stability. The study promotes a shift toward stability-aware forecasting practices, offering practical guidelines for building more robust and sustainable prediction systems.
keywords:
Time series, Demand forecasting, Forecasting competitions, Cross-learning, Global models, Forecast stability , Vertical stability , Machine learning, Deep learning, Conformal predictions
1 Introduction
In recent years, global forecasting models, those trained across multiple time series simultaneously, have emerged as a powerful alternative to traditional local approaches, particularly in large-scale applications such as retail demand forecasting. While their accuracy and efficiency have been widely studied, less attention has been paid to the stability of their forecasts over time. In many operational contexts, forecasts are produced regularly and serve as the foundation for critical business decisions, from inventory planning to resource allocation. In such settings, forecast stability becomes a key requirement; not only should forecasts be accurate, but they should also remain reasonably consistent over time. Indeed, forecast stability refers to the consistency of predictions produced by a forecasting model as new data becomes available. As defined by Godahewa et al. (2025), forecast stability can be categorized into two primary forms: vertical stability, which concerns the consistency of forecasts for the same target date across different forecast origins, and horizontal stability, which addresses the smoothness of forecasts across the forecast horizon from a single origin. Both types play a distinct role: vertical stability helps avoid costly forecast revisions that can disrupt planning cycles, while horizontal stability prevents erratic behavior across time steps that could lead to inefficient operations and amplification of demand fluctuations, such as the bullwhip effect in supply chains (Lee et al., 1997). In particular, vertical stability may also be seen as temporal robustness, in the sense that a forecasting model is robust to updates (or retraining) as new observations are available, a property that is critical in practice for avoiding frequent changes in business decisions. Forecast instability, indeed, can have serious consequences. Unstable forecasts may lead to frequent and expensive adjustments in supply chain plans, diminish trust in the forecasting system, and complicating the decision-making processes, leading to suboptimal business outcomes.
Nevertheless, despite the importance of forecast stability, it is still common practice to evaluate forecasting models only on their accuracy, even in frequent retraining settings where stability is more likely to be compromised. This can also be due to an inherent perception that there exists a trade-off between accuracy and stability. As new information becomes available, updated forecasts may naturally differ from earlier versions, ideally improving in accuracy at the expense of stability. Conversely, overly stable forecasts might sacrifice accuracy by ignoring valuable new information. Ensembling techniques, which integrate multiple forecasts from different models, have been proposed as a mechanism to concurrently reduce bias and variance while enhancing stability (Wang et al., 2023). By aggregating diverse predictions, ensembling methods can alleviate the impact of individual model volatility and potentially achieve a more favorable equilibrium between accuracy and stability. Even if the trade-off between forecasting accuracy and stability presents a significant challenge for organizations seeking to balance predictive performance with operational consistency, recent research suggests that stability and accuracy are not necessarily in conflict, and both can be achieved at the same time ((Godahewa et al., 2025), (Van Belle et al., 2023)). Moreover, Zanotti (2025b) demonstrated that the retraining frequency does not harm the forecast accuracy of global models, meaning that models estimated using this forecasting approach retain the same level of performance even when updated less frequently. This has a strong potential implication on the forecast stability: if we can reduce the retraining frequency of global models, then we should obtain more stable predictions, because stability should be a non-increasing function of the retraining frequency.
1.1 Research Question
We aim to address the question ”Are global forecasting models stable?”. Purposely, we study whether the global modeling approach produces stable forecasts, and we try to understand the effects of retraining on the forecast stability, that is whether avoiding re-estimation for every new observation damages the stability of global models. To address this question, we rely on the two most recent and comprehensive retail forecasting datasets: the M5 and VN1 competition data.
To generally understand the stability (or instability) of the forecasting models, we consider ten distinct global forecasting methods (five from the traditional machine learning domain and five based on commonly used deep neural network architectures). Moreover, we examine stability across a range of retraining scenarios, from no retraining to continuous retraining, by exploring periodic strategies that broadly encompass the most practical and effective approaches.
We also investigate the use of ensembling, or forecasting combinations, as a mean to obtain more stable forecasts. This approach can indeed be significant to mitigate forecast instability, reducing the model variance and bias.
1.2 Contributions
Our contribution is fourfold:
  • 1.
    We provide the first comprehensive study of the forecast stability of global models, using 10 distinct methods, a diverse collection of real-world datasets, and evaluating both point and probabilistic predictions.
  • 2.
    We analyze the relationship between retraining frequency and forecast stability, comparing different scenarios to quantify the impact of frequent retraining in terms of the stability of forecasting.
  • 3.
    We suggest a new metric to evaluate the instability of probabilistic forecasts.
  • 4.
    We present practical guidelines for organizations and practitioners on when and how often to retrain global forecasting models to obtain stable forecasts.
By tackling these aspects, this paper contributes to both the forecasting and machine learning communities by providing insights into the stability of global forecasting models.
1.3 Overview
The rest of this paper is organized as follows. After a brief review of related works (Section 2), in Section 3 we describe the design of the experiment used in our study. The datasets and their characteristics are presented in 3.1, and the methods adopted for global forecasting are discussed in 3.2. The concepts related to model update and retrain scenario are explained in 3.3, together with the evaluation strategy adopted, while the metrics used to assess the stability of models are shown in 3.4. In Section 4 we discuss the empirical findings of our study, including forecast stability and ensembling on the different scenarios. Finally, Section 5 contains our summary and conclusions.
2 Related works
The cross-learning approach has seen substantial development in recent years. Today, most time series forecasting studies include at least a benchmark comparison involving global models (GMs), underscoring their growing importance in the field. Semenoglou et al. (2021) demonstrated the high accuracy of GMs on the M4 competition dataset, while Hewamalage et al. (2022) explored the conditions under which global models are competitive. Additionally, Montero-Manso & Hyndman (2021) and (Montero-Manso, 2023) provided theoretical support showing that GMs can match or surpass the accuracy of local models, with lower complexity and without assuming data similarity. Global models have proven to be the most accurate method in several forecasting domains, including retail demand (Spiliotis et al. (2022), Bandara et al. (2019), Juan R Trapero & Fildes (2015)), electricity demand (Buonanno et al., 2022), water demand (Groß & Hans, 2024), gas consumption (Gaweł & Paliński, 2024), and crop production (Ibañez & Monterola, 2023). Their effectiveness was particularly evident during the M5 competition (Makridakis et al., 2022a), where tree-based models employing cross-learning ranked among the top-performing solutions (Januschowski et al., 2022). To further enhance GM performance, several strategies such as clustering (Godahewa et al. (2021a), Bandara et al. (2020)) and data augmentation (Bandara et al., 2021) have been explored. Moreover, new machine learning (Godahewa et al., 2023) and deep learning (Oreshkin et al., 2020) architectures have been specifically designed to support cross-learning. Recently, research has begun focusing on improving GMs’ ability to capture local patterns (Sen et al., 2019) and enhance their interpretability (Rajapaksha et al., 2023).
From a forecasting stability perspective, the literature is very poor. Godahewa et al. (2025) introduced the first classification of the different types of forecasting stability, that is horizontal or vertical stability, proposing a new model-agnostic framework based on linear combinations of predictions to obtain more stable forecasts. Van Belle et al. (2023), instead, extended an existing deep learning architecture (NBEATS) to optimize forecasts from both a traditional forecast accuracy perspective as well as a forecast stability perspective, directly including an instability component into the loss function of the model. There is also active research on forecast stability within judgmental forecasting ground (Fildes & Goodwin, 2021). However, almost all the literature related to forecast stability is focused on point prediction stability only, possibly because it is not clear how to measure instability in terms of probabilistic forecasting, and maybe also because most machine learning and deep learning methods do not directly output probabilistic predictions (Makridakis et al., 2022c). Nevertheless, in many forecasting applications, such as supply chain management, it is crucial to generate and evaluate predictions probabilistically, whether through prediction intervals, quantiles, or full predictive distributions (Fildes et al., 2022). Among the methods developed for uncertainty quantification, Vovk et al. (2005) introduced Conformal Inference, a model-agnostic framework that offers valid uncertainty estimates and can also be applied in time series forecasting settings (Stankeviciute et al., 2021).
In the context of model retraining and updating strategies, the most comprehensive work in time series forecasting is by Zanotti (2025b), who showed the positive effects of reducing the retraining frequency of global models on both accuracy and forecasting costs. Spiliotis & Petropoulos (2024) extensively investigated the impact of various retraining scenarios and parameter update methods on model performance. However, their focus was limited to the exponential smoothing family, following the traditional local modeling approach. (Huber & Stuckenschmidt, 2020) touched on retraining within the retail demand domain, but the study was restricted to a small set of models, limited retraining strategies, and a proprietary daily dataset. While these findings are encouraging, there has been little direct investigation into whether global models specifically can retain stability with less frequent updates.
Building on these prior studies, our work directly examines the stability of global models. Moreover, by systematically evaluating a wide range of retraining strategies and their effects on the forecasting stability of various global models, we seek to offer both theoretical insights and practical guidance for promoting more stable forecasting practices.
3 Experimental design
This section presents the empirical analysis carried out to investigate the stability of global models and to determine whether less frequent retraining scenarios can yield stability outcomes comparable to those of the baseline scenario, which involves the most frequent retraining. We begin by describing the datasets used in our experiments, followed by an overview of the machine learning, deep learning, and ensemble models employed. Finally, we detail the instability measures, the various retraining scenarios considered, and the evaluation strategy applied to assess forecast performance.
3.1 Datasets
For our experiments, we employed two retail forecasting datasets: the M5 and the VN1 competition datasets. The M5 competition, part of the well-known M-competitions series organized by Spyros Makridakis and colleagues, aimed to benchmark forecasting methods in the context of retail demand (Makridakis et al., 2022b). The M5 dataset (Howard & Makridakis, 2020) includes 3,049 daily time series representing unit sales of Walmart products across three categories (Food, Hobbies, and Household) sold in ten stores located in California, Texas, and Wisconsin. The data span from 2011 to 2016 and are characterized by high intermittency and a hierarchical structure, enabling forecasts at multiple aggregation levels (e.g., SKU, product category, store, and state). The dataset also includes exogenous variables such as prices, promotions, and special events (e.g., holidays), which can influence demand. The VN1 Forecasting - Accuracy Challenge, organized in October 2024 by Flieber, Syrup Tech, and SupChains, represents the first edition of a new competition series (Vandeput, 2024). The dataset comprises weekly sales data for 15,053 products sold by U.S.-based e-vendors between 2020 and 2024. Unlike the M5, which involves a single retailer (Walmart) and a limited number of physical stores, the VN1 dataset captures sales from 328 warehouses operated by 46 different retailers. As far as we are aware, our study is among the first to evaluate forecasting models on this dataset. Together, these two datasets offer the most recent and comprehensive time series collections related to retail demand, enabling good generalizability of our findings.
Table 1: The M5 and the VN1 datasets used in the experiments.
Dataset Frequency N. Series Min Obs per Series
M5 Daily (7) 28.298 730
VN1 Weekly (52) 15.053 157
In both cases, our analysis focused on the most disaggregated level of the data (i.e., SKUs), where the potential advantages of reducing retraining frequency are likely to be most pronounced. To ensure consistency with the evaluation setup described in Section 3.3, we applied a filtering criterion: for the daily data (M5), only time series with at least two years of observations (730 data points) were retained; for the weekly data (VN1), we considered only those series with at least three years of data (157 observations).
3.2 Forecasting models
In this section, we present an overview of the global models used in our experiments.
Following Zanotti (2025b), let us define $\mathcal{Y}$ as the set of all available time series in a dataset, where each $Y_{i}\in\mathcal{Y}$ represents an individual time series. Let $\mathcal{F}$ denote the set of possible predictive functions, with each $F\in\mathcal{F}$ corresponding to a specific forecasting model. Without loss of generality, we assume that all necessary information for prediction is contained within $\mathcal{Y}$. Under the local modeling paradigm, forecasts for a horizon $h$ are generated by training a separate model for each individual time series. This implies that each series $Y_{i}$ is associated with its own model, characterized by its own set of parameter values.
$Y_{i}^{h}=F(Y_{i},\theta_{i})\mbox{.}$ (1)
In contrast, under the global modeling framework, forecasts for each individual time series are generated using a single model trained on the entire dataset. This approach leverages cross-series information, allowing the model to learn shared patterns and structures across all time series in $\mathcal{Y}$.
$Y_{i}^{h}=F(\mathcal{Y},\Theta)\mbox{.}$ (2)
It is important to note that in the cross-learning methodology, the model parameters $\Theta$ are not specific to individual time series but are shared across all series in the dataset. This parameter sharing is a key characteristic of global models, enabling them to generalize patterns across series and potentially reducing the forecast instability.
In our study, we focused exclusively on analyzing the stability of global methods, as our primary goal was to assess whether this modeling approach, unlike the traditional local one, can maintain performance with less frequent retraining. Cross-learning has become the standard in many industries dealing with large-scale time series data, particularly in retail demand forecasting, where regular forecasts for thousands of SKUs are required (Januschowski et al., 2020). Therefore, to conduct a comprehensive evaluation of the stability of global models, we incorporated both traditional machine learning algorithms and state-of-the-art deep learning architectures. The selected models are well-established in the time series forecasting literature and represent a range of methodological paradigms, allowing for a broad and informative comparison.
As in Zanotti (2025b), we used five machine learning models and five deep learning models. Machine learning models have proven effective in forecasting tasks due to their ability to capture complex non-linear relationships in data. They are also relatively easy to train and tend to deliver strong performance when leveraging cross-learning techniques. In this study, we experimented with Linear (Pooled) Regression and four widely used tree-based methods. While machine learning models are easier to train compared to deep learning alternatives, they often rely heavily on high-quality feature engineering (Januschowski et al., 2022). For our experiments, we adopted simplified versions of the feature engineering pipelines used in the top-performing M5 and VN1 solutions. We constructed time series features such as lags, rolling means, expanding means, and calendar features (e.g., year, month, week, day of the week). We also included static metadata such as store, product, category, and location identifiers based on the dataset’s frequency. For the M5 dataset, additional external features like special events were also incorporated. Model hyperparameters were selected based on configurations used in top-performing competition solutions when available; otherwise, we relied on the recommended defaults from the respective libraries. We compared five machine learning models for time series forecasting. Linear Regression (LR) is a classical statistical model effective with appropriate feature engineering, often used as a benchmark in global model performance ((Montero-Manso & Hyndman, 2021), (Godahewa et al., 2021b)). Random Forest (RF), an ensemble learning method, captures non-linear patterns and is robust to overfitting, making it suitable for demand forecasting ((Breiman, 2001), (Januschowski et al., 2022)). Extreme Gradient Boosting (XGBoost) and Light Gradient Boosting Machine (LGBM) are gradient-boosted models known for speed and accuracy, with LGBM excelling in large, high-dimensional datasets ((Chen & Guestrin, 2016), (Ke et al., 2017), Makridakis et al. (2022a)). Categorical Boosting (CatBoost) specializes in handling categorical data with minimal preprocessing (Prokhorenkova et al., 2018).
Deep learning models have gained significant traction in time series forecasting due to their ability to model long-range temporal dependencies and to learn hierarchical representations directly from raw input data (Goodfellow et al., 2016). Unlike machine learning models, deep learning architectures typically do not require extensive manual feature engineering. They can autonomously learn lagged structures, trend, and seasonality directly from the raw series. Nevertheless, training deep models is often more complex due to the higher number of hyperparameters and their sensitivity to configuration choices, which can significantly affect performance (Smyl, 2020). In our implementation, we followed competition-proven practices by providing only static, calendar, and external covariates as inputs, while using top solutions’ guidelines for setting the model hyperparameters. We compared five deep learning models for time series forecasting. Multi-Layer Perceptron (MLP) is a versatile, efficient neural network used in many time series forecasting tasks (Rosenblatt, 1958), while Recurrent Neural Networks (RNNs), particularly LSTMs and GRUs, capture temporal dependencies in sequential data ((Cho et al., 2014), (Hochreiter & Schmidhuber, 1997)). Temporal Convolutional Networks (TCN) offer an alternative by using causal convolutions to capture long-range dependencies (Van den Oord et al., 2016). Neural Basis Expansion Analysis for Time Series (NBEATS) and Neural Hierarchical Interpolation for Time Series (NHITS) are state-of-the-art models for time series forecasting, with NBEATS offering interpretable trend and seasonality components (Oreshkin et al., 2020), and NHITS improving upon NBEATS with hierarchical interpolation mechanisms (Challu et al., 2022).
Moreover, we also estimated four different ensemble learning models based on Zanotti (2025a). Ensemble learning combines the predictions of multiple base models to improve forecast accuracy and robustness, particularly when individual models capture complementary patterns in the data. In the context of global forecasting, ensembles are especially valuable for mitigating the risk of model instability and overfitting, which can occur when relying on a single method (Wang et al., 2023). In our study, we adopted a simple averaging strategy, which is widely used for its ease of implementation and effectiveness in improving forecast robustness (Claeskens et al., 2016). Specifically, we constructed ensembles by averaging the forecasts of the top two, top three, top four, and top five most accurate individual models, as determined on out-of-sample performance. The ensemble configurations were designed to reflect practical setups and were validated using the same rolling origin evaluation framework described in Section 3.3. By incorporating ensemble learning into our analysis, we aimed to assess whether combining global models could further enhance forecasting stability under different retraining scenarios. Moreover, by evaluating multiple ensemble sizes, we aimed to understand how the inclusion of additional models affects the stability under different circumstances.
All global models were implemented in Python using Nixtla’s framework (Nixtla, 2022). Specifically, the mlforecast library was used to train the machine learning models, while neuralforecast was employed for efficiently training the deep learning models.
3.3 Evaluation strategy
Out-of-sample evaluation is a cornerstone of time series forecasting, providing a way to test a model’s generalization ability beyond the training period, an essential step given that future data may diverge from historical patterns due to structural breaks, level shifts, or unexpected shocks (Tashman, 2000). Among the various evaluation strategies, the rolling origin evaluation has emerged as the most widely accepted and rigorous method (Bergmeir & Benítez, 2012). It respects the chronological nature of time series data while enabling repeated assessments across multiple forecast origins. In this framework, the time series is split into a training set and a test set, where the model is trained on the former and evaluated on the latter. Forecasts are generated for a defined horizon $h$, and at each iteration, the forecast origin is shifted forward by a specified step size, typically one, to simulate real-time forecasting. The model is then retrained on the updated training data, using either a fixed-length or expanding window. Stability metrics (see Section 3.4) are averaged over all iterations to provide a robust estimate of forecasting accuracy.
Compared to fixed origin evaluations, rolling origin offers a more nuanced view of a model’s robustness by exposing it to a range of temporal conditions (including seasonal patterns, trends, and potential anomalies) that may not be captured in a single evaluation window (Bergmeir & Benítez, 2012). This iterative approach reduces the risk of overfitting to one specific train-test split, and is especially beneficial in operational contexts like retail, logistics, or finance, where forecasts are regularly updated in response to incoming data. It also aligns well with real-world decision-making processes that require forecasts to adapt over time.
In practice, the design of a rolling origin evaluation depends on several parameters: the size of the test set, the length of the forecast horizon, the step size, and the windowing strategy (fixed vs expanding). The expanding window strategy is particularly suited for short time series, as it maximizes the amount of historical data available at each iteration. This approach is widely used in applied settings (Petropoulos & et al., 2022), and it is the one we adopted in our study, especially given the relatively short length of the weekly VN1 series. Following Zanotti (2025b), Table 2 outlines the key parameters used in our experiments, including a test set spanning at least one full year to mitigate seasonal bias, horizons aligned with typical business use cases, and a step size of one to maximize evaluation frequency.
Table 2: The retraining scenarios, the test window, and the horizon for the M5 and the VN1 datasets.
Dataset Frequency Retraining Scenarios (r) Test Window (T) Horizon (h)
M5 Daily (7) 7, 14, 21, 30, 60, 90, 120, 150, 180, 364 364 28
VN1 Weekly (52) 1, 2, 3, 4, 6, 8, 10, 13, 26, 52 52 13
Furthermore, to investigate the impact of retraining frequency on forecasting stability, we examined various retraining scenarios, or retraining windows, as defined in Zanotti (2025b). Each window $r$ indicates the number of new observations after which the model is retrained. These retraining scenarios are tailored to the data frequency, daily for M5 and weekly for VN1, since frequency dictates both the forecast horizon and business review cycles. Table 2 summarizes the selected retraining windows, test periods, and forecast horizons. For instance, in the daily case, $r=7$ reflects weekly updates. Across all scenarios, we trained global models on aligned datasets, updating the model either completely at each retraining step or not at all. We excluded hyperparameter tuning due to its high cost and marginal expected benefit. Following Zanotti (2025b), we treated $r=1$ (or $r=7$ for daily data) as the accuracy benchmark and $r=T$ as the no-retraining baseline, with intermediate values representing periodic retraining strategies.
3.4 Evaluation metrics
3.4.1 Stability metrics
The stability evaluation of point predictive models is a crucial topic in time series forecasting: very few metrics are available to capture models’ vertical stability, and there is no consensus in the literature on what metric to use (Godahewa et al., 2025). Van Belle et al. (2023) proposed the symmetric Mean Absolute Percentage Change (sMAPC), which measures the change of one to $h$-step ahead forecasts obtained by two consecutive forecast origins, providing a measurement of up to which extent the forecasts generated at the first origin are unstable compared to those generated at the second origin.
$\text{sMAPC}=\frac{200}{h-1}\sum_{t=n+1}^{n+h-1}\frac{\lvert\hat{y}_{t,n}-\hat{y}_{t,n-1}\rvert}{\lvert\hat{y}_{t,n}\rvert-\lvert\hat{y}_{t,n-1}\rvert}\mbox{.}$ (3)
Here, $\hat{y}_{t,n}$ and $\hat{y}_{t,n-1}$ are the forecasts generated for period $t$ with origins $n$ and $n-1$ respectively. The instability across different pairs of consecutive forecasting origins can be obtained by a simple average of the sMAPC values. Lower values imply less unstable predictions.
Beyond point forecasts, our study also emphasizes the probabilistic stability of the models under different retraining strategies. Since most machine learning and deep learning models do not natively produce full predictive distributions, we employed Conformal Inference to construct prediction intervals around point forecasts. Conformal Inference is a flexible and robust framework that provides distribution-free, model-agnostic prediction intervals with guaranteed finite-sample coverage under mild assumptions (Vovk et al., 2005). Although initially developed for i.i.d. data, recent extensions have adapted it for time series applications by relaxing the exchangeability assumption (Stankeviciute et al., 2021). Its advantages (validity guarantees, minimal assumptions, computational simplicity, and effectiveness with limited data) make it especially suitable for our global forecasting setup, where different models are compared across heterogeneous datasets and varying levels of retraining frequency. Because no metric exists to evaluate the vertical stability (or instability) of the resulting probabilistic forecasts, we proposed a new measure, the Multi-Quantile Change (MQC, or Multi-Quantile Loss Change, MQLC), to comprehensively assess the stability of the probabilistic predictions. We defined the Quantile Change and the Multi-Quantile Change as:
$\text{QC}=\frac{1}{h-1}\sum_{t=n+1}^{n+h-1}\left(q\cdot(\hat{y}_{t,n-1}-\hat{y}_{t,n})\cdot\mathbb{I}_{\hat{y}_{t,n-1}\geq\hat{y}_{t,n}}+(1-q)\cdot(\hat{y}_{t,n}-\hat{y}_{t,n-1})\cdot\mathbb{I}_{\hat{y}_{t,n-1}<\hat{y}_{t,n}}\right)\mbox{,}$ (4)
$\text{MQC}=\frac{1}{\mathcal{Q}}\sum_{q\in\mathcal{Q}}\text{QC}(q)\mbox{.}$ (5)
The Quantile Change (QC) metric is a measure of the change in forecasted quantiles between two consecutive forecast origins. Given two forecast origins, $n$ and $n-1$, QC quantifies the average adjustment in the predicted quantiles across the forecast horizon $h$. The formula mirrors the Quantile Loss (Pinball Loss) but replaces the actual observations $y_{t}$, with previous forecasts $\hat{y}_{t,n-1}$, thereby shifting the focus from forecast accuracy to forecast stability. The intuition is straightforward: if the forecasts are stable across time (i.e., $\hat{y}_{t,n}\approx\hat{y}_{t,n-1}$), then QC will be small. Large QC values indicate that the model updates its distribution significantly with each new observation, suggesting a lack of temporal robustness in its probabilistic output. By aggregating this across a range of quantiles $\mathcal{Q}$, the Multi-Quantile Change (MQC) captures overall instability across the entire forecast distribution, providing a comprehensive metric for probabilistic stability. One drawback of the MQC is that it is based solely on a finite set of quantiles rather than the full predictive distribution. As a result, it captures changes only at selected points in the distribution (e.g., median, tails), potentially missing more subtle shifts in shape, skewness, or variance that occur between those quantiles. However, in practical forecasting applications, especially in retail and supply chain domains where decisions are often made based on specific quantile levels (e.g., high quantiles for safety stock), this limitation is mitigated by the fact that the selected quantiles are usually those of greatest operational interest. Moreover, through Conformal Inference, it is possible to calculate as many quantiles as desired at no computational cost, making MQC a pragmatic and model-agnostic metric for capturing distributional instability across time.
We computed 13 quantiles: the median and 6 central prediction intervals (60%, 70%, 80%, 90%, 95% and 99%). Central intervals describe the forecast center, while wider intervals capture tail risks, essential for safety stock decisions in retail (Barrow & Kourentzes, 2016). To ensure reliable quantile estimates, conformal prediction intervals were computed on a validation set at least twice the forecast horizon, which constrained the number of time series used.
We normalized each evaluation metric relative to the baseline retraining scenario, defined by the dataset frequency, to enable consistent comparison across models and retraining windows. To statistically validate our findings, we applied the Friedman-Nemenyi test for multiple comparisons (Demšar, 2006).
3.4.2 Performance metrics
Evaluating point forecast accuracy in time series is a debated topic, with no consensus on the best metric (Hewamalage et al., 2023). In our context of SKU-level demand forecasting, characterized by intermittent data, metrics based on absolute or percentage errors are suboptimal (Kolassa, 2020), and due to varying scales across series, a scaled accuracy metric has to be preferred. Hence, we adopted the Root Mean Squared Scaled Error (RMSSE) (Hyndman & Koehler, 2006),
$\text{RMSSE}=\sqrt{\frac{\frac{1}{h}\sum_{t=n+1}^{n+h}(y_{t}-\hat{y}_{t})^{2}}{\frac{1}{n-s}\sum_{t=s+1}^{n}(y_{t}-y_{t-s})^{2}}}\mbox{.}$ (6)
which compares the model’s squared error to that of a seasonal naive forecast, and it was the official accuracy metric in the M5 competition (Makridakis et al., 2022b). Lower values indicate better performance.
To assess probabilistic accuracy, we evaluated the quantiles produced through the Conformal Inference framework using the Quantile Loss (QL) and Multi-Quantile Loss (MQL).
$\text{QL}=\frac{1}{h}\sum_{t=n+1}^{n+h}\left(q\cdot(y_{t}-\hat{y}_{t})\cdot\mathbb{I}_{y_{t}\geq\hat{y}_{t}}+(1-q)\cdot(\hat{y}_{t}-y_{t})\cdot\mathbb{I}_{y_{t}<\hat{y}_{t}}\right)\mbox{,}$ (7)
$\text{MQL}=\frac{1}{\mathcal{Q}}\sum_{q\in\mathcal{Q}}\text{QL}(q)\mbox{.}$ (8)
These proper scoring rules assess forecast distribution accuracy (Kolassa, 2016). Moreover, the MQL was the official metric of the M5 Uncertainty competition (Makridakis et al., 2022c).
For this study, we adopted a cloud computing machine NC6s_v3 hosted on Microsoft Azure, with Linux Ubuntu 24 operating system, 1 Graphical Processing Unit, 6 Computing Processing Units, 112GB of memory.
4 Results and discussion
This section discusses the empirical findings of our study, highlighting the interplay between accuracy, probabilistic performance, and stability cost across different retraining strategies and ensemble configurations. We extract insights from both the M5 and VN1 datasets to uncover consistent patterns as well as nuances unique to each dataset’s characteristics.
Table 3: Overall forecasting performance and stability for each method across datasets. RMSSE, MQL, sMAPC, and MQC. Minimum values per column are highlighted in bold.
Method M5 VN1
RMSSE MQL sMAPC MQC RMSSE MQL sMAPC MQC
LR 0.777 0.267 0.188 0.118 6.549 2.896 0.085 1.064
RF 1.868 2.590 0.092 1.322
XGBoost 0.755 0.258 0.075 0.115 1.890 2.469 0.146 1.101
LGBM 0.771 0.256 0.054 0.123 3.542 2.625 0.121 1.282
CatBoost 0.947 0.263 0.118 0.127 5.762 2.845 0.320 1.435
MLP 0.821 0.281 0.500 0.109 1.543 2.492 0.458 0.844
LSTM 1.913 2.843 0.558 1.020
TCN 0.865 0.290 0.499 0.114 1.913 2.843 0.600 1.020
NBEATSx 0.815 0.279 0.547 0.110 1.698 2.626 0.596 1.532
NHITS 0.828 0.284 0.566 0.112 1.699 2.632 0.605 1.592
Ens2A 0.757 0.255 0.059 0.117 1.472 2.369 0.503 1.079
Ens3A 0.757 0.256 0.089 0.117 1.517 2.410 0.514 1.200
Ens4A 0.758 0.249 0.095 0.113 1.524 2.386 0.167 1.180
Ens5A 0.763 0.251 0.099 0.111 1.544 2.375 0.164 1.153
Table 3 summarizes the performance and stability of all forecasting methods evaluated in this study, across both the M5 and VN1 datasets. For each method, the table reports four key metrics per dataset: RMSSE (Root Mean Squared Scaled Error) to evaluate point forecast accuracy, MQL (Multi-Quantile Loss) to assess probabilistic forecast performance, sMAPC (symmetric Mean Absolute Percentage Change) to compare the stability in point forecast settings, and MQC (Multi-Quantile Change) to understand the stability in probabilistic terms. Overall 111 The overall results are based on the baseline retraining scenario, $r=7$ for M5 and $r=1$ for VN1, which is commonly considered the standard in both theoretical literature and practical forecasting applications. , the models evaluated achieved better stability results on the M5 dataset than on VN1. Several factors may account for this difference, such as the larger dataset size and higher frequency of the M5 time series, the availability of rich external regressors (e.g., promotions, special events), and the existence of well-established benchmark hyperparameter settings, many of which were unavailable for VN1 at the time of model training. Moreover, in terms of point forecasting, machine learning models are much more stable than deep learning methods on both datasets, implying that the former are less sensitive to frequent updates. On the contrary, in terms of probabilistic forecasting, deep learning architectures show less instability, meaning that these approaches can be a better option when retraining is performed frequently.
In general, ensembling seems to be a good solution to produce stable predictions. However, combining the most accurate models is often not sufficient to reduce the instability. Indeed, ensembles show increasing returns in stability as more models are added, as opposed to the well-known phenomenon in the ensemble literature, where accuracy improvements show diminishing returns from adding more models (Zanotti, 2025a). Interestingly, larger ensembles are usually more effective for stability purposes, implying that model diversity may be a more relevant criterion to create a forecast combinations, in contrast to model performance. This pattern is particularly true for probabilistic forecasting, but it is also evident in point predictions. For instance, in the VN1 dataset, the Ens4A and Ens5A combinations achieve good stability results simply because they contain both deep learning and machine learning models, extending the combination’s diversity. Indeed, even if the deep learning models’ instability is on average 60%, it is enough to add one single machine learning model to obtain an improvement of almost 45%.
Figure 1 shows the point forecast stability of each model along the different retraining scenarios for the M5 and the VN1 datasets. To allow comparisons, we display the results in relative terms with respect to the baseline scenarios, that is $r=7$ for M5 and $r=1$ for VN1. As expected, the sMAPC profiles are non-increasing functions of the retraining frequencies. The stability remains practically the same across update frequencies or even improves as the retraining period increases. Indeed, moving from high to low retraining frequencies, the stability of most global models improves compared to the baseline. In particular, for some models (i.e. CatBoost, LSTM, and TCN) the instability reduction is strong and consistent, reaching very low levels. These results imply that less frequent retraining may have a huge impact on the point forecast stability of global models. This can be explained by the fact that, when a forecasting origin is updated and the model is retrained on the new data, the model’s predictions may differ consistently (even using the cross-learning approach) from those produced before the update. Therefore, less retraining is often synonymous with more stability.
Figure 1: sMAPC results for each method and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
Figures 2, and 3 statistically confirms the above results for the M5 and VN1 datasets. It is clear that most periodic retraining scenarios are statistically different to the continuous retraining strategy in terms of point forecast stability at the 5% level.
Figure 2: M5 Friedman-Nemenyi test results based on sMAPC.
Figure 3: VN1 Friedman-Nemenyi test results based on sMAPC.
In a similar fashion, Figure 4 summarizes the relative stability in a probabilistic forecasting setting. In this context, we observe that, for the M5 dataset, the stability (as measured by Multi Quantile Change) paths are constant across almost every retraining scenario, meaning that less frequent updates neither harm nor improve the probabilistic forecasting stability of most global models. For the VN1 dataset, instead, we observe an almost convex relationship between the stability and the retraining frequency. On average, models’ stability improves for low retraining scenarios and then starts deteriorating around $r=10$, implying that it is possible to avoid retraining up to 10 weeks without negatively impacting on the forecast stability. For both datasets, these results are confirmed by the statistical tests on MQC over the different retrain periods (see supplementary materials).
Figure 4: MQC results for each method and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
Overall, the results from Figures 1 and 4, supported by the Friedman-Nemenyi tests, indicate that reducing the retraining frequency of global models does not negatively impact (and may even enhance) forecast stability, for both point and probabilistic forecasts. When considered alongside the findings of Zanotti (2025b), this provides strong evidence against the practice of continuous retraining in global forecasting models. Lower retraining frequencies preserve both accuracy and stability, while offering substantial savings in computational resources.
Figure 5 shows the evolution of point forecast stability, across different retraining scenarios for various ensemble configurations. The results on the M5 dataset indicate that reducing retraining frequency may negatively affect forecast stability. However, this impact is below 1% for most retraining scenarios. Indeed, statistical tests (in supplementary material) on the significance of these differences show that most retraining scenarios are not statistically different from each other, with periodic retraining strategies being at least as good as continuous retraining. The results on the VN1, instead, indicate that there exists a clear benefit in lowering the retraining frequency in terms of forecast stability. Notably, on the M5 dataset the smallest ensemble (Ens2A) consistently achieves the lowest sMAPC values, since it is obtained by combining XGBoost and LGBM that where both the most accurate and the most stable global forecasting models on that set of data. Nonetheless, larger ensembles (such as Ens4A and Ens5A) stability profiles tend to outperform smaller ones, especially at lower retraining frequencies, suggesting that model diversity plays a critical role in stabilizing ensemble forecasts when. Figure 6 further supports these observations in the probabilistic domain. First of all, probabilistic forecasts of ensembles are generally more stable than that of the base models across retraining frequencies, especially in M5, where MQC variations are within 1 percentage point. Moreover, in the VN1 dataset, the almost convex relationship between stability and the retraining frequency is flattened, because the variations are more restrained as the retraining period increases, and the decline in stability happens only from $r=26$ onwards. Then, larger ensembles provide consistently more stable probabilistic performance across most retraining scenarios, confirming the role of the diversity composition of ensembles even in the probabilistic setting. Most importantly, the effect of ensembling on the profiles of stability, as measured by sMAPC or MQC, is that it greatly reduces the variations across the different retraining frequencies, reducing both positive and negative stability changes.
Figure 5: sMAPC results for each ensemble and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
Figure 6: MQC results for each ensemble and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
Figures 7 and 8 illustrate the fundamental trade-off between forecast accuracy and stability, comparing RMSSE with sMAPC and MQL with MQC, respectively. These plots reveal that high accuracy does not guarantee high stability. In particular, several deep learning models, achieve low RMSSE or MQL scores but exhibit high instability, especially in VN1, highlighting their sensitivity to changes in the training data and their volatility across forecast origins. Conversely, models like Random Forest, XGBoost, and LGBM strike a more favorable balance, maintaining relatively low RMSSE and sMAPC values. Nevertheless, it is the ensemble models, and especially those with larger size like Ens4A and Ens5A, that most consistently approach the optimality of stability and accuracy. In both the point and probabilistic domains, even very simple ensembles manage to control variance and mitigate instability without sacrificing much, if any, predictive performance. The overall picture that emerges from these figures reinforces the idea ensembling offers a practical and effective strategy to counteract forecast instability, particularly when constructed from a diverse pool of models.
Figure 7: The accuracy-stability trade-off. RMSSE vs sMAPC for the M5 and VN1 datasets.
Figure 8: The accuracy-stability trade-off. MQL vs MQC for the M5 and VN1 datasets.
This set of results has clear practical implications: in operational forecasting systems, especially in domains such as retail and supply chain management, the use of ensembles combined with lower retraining frequencies may offer a robust forecasting process, balancing performance with reliability.
5 Conclusions
This study provides a comprehensive empirical assessment of forecast stability in global models, focusing on the effects of retraining frequency and ensembling across both point and probabilistic forecasting contexts. Our findings demonstrate that global forecasting models are able to produce very stable predictions and that they can achieve even more stability under reduced retraining frequencies, challenging the prevailing wisdom that frequent model updates are necessary to maintain forecast reliability.
In terms of point forecasting, the results consistently show that retraining less frequently does not harm, and often improves, forecast stability. This trend is particularly evident in the M5 dataset, where many models exhibit monotonic or flat stability profiles as the retraining interval increases. Similar patterns, although more nuanced, are observed in the VN1 dataset. Probabilistic forecast stability, as measured by the newly defined Multi-Quantile Change (MQC), shows a slightly different behavior. While in the M5 dataset forecast stability remains largely unaffected by retraining frequency, in VN1 we observe a convex relationship, indicating that retraining up to every 10 weeks can preserve stability, but beyond that point the quality of probabilistic forecasts begins to deteriorate. These results are statistically validated via Friedman-Nemenyi tests and indicate that continuous retraining is often not the optimal strategy.
Ensembling emerges as a powerful strategy to enhance forecast stability. While combining only the most accurate models offers moderate gains, the most substantial improvements in stability, both point and probabilistic, are achieved by ensembles that incorporate a diverse mix of forecasting algorithms. This effect is especially pronounced in the VN1 dataset, where larger and more heterogeneous ensembles significantly reduce forecast volatility across retraining scenarios. Notably, ensemble models not only improve stability metrics but also smooth the variation in stability across retraining frequencies, further reinforcing their role as a stabilizing force in global forecasting systems. These results suggest that model diversity should be prioritized over raw model performance when building ensembles.
The joint analysis of retraining and ensembling reveals that the two approaches act as complementary levers for stabilizing forecasts. Less frequent retraining reduces abrupt changes in model behavior across forecasting origins, while ensembling mitigates the instability arising from any single model’s sensitivity to input changes. When combined, these strategies offer a robust forecasting setup that minimizes the instability of both point estimates and prediction intervals without sacrificing predictive accuracy. Figures comparing the accuracy–stability trade-off clearly show that ensembles lie closer to the optimality, offering an optimal balance between these competing objectives.
From a practical standpoint, these findings have several implications. First, businesses that rely on large-scale forecasting systems, such as those in retail, inventory management, and supply chain planning, can adopt lower retraining frequencies not only to conserve computational resources but also to increase operational stability, without compromising on forecast quality. Second, incorporating diverse ensembles into global forecasting frameworks provides an effective hedge against model instability, thereby increasing the reliability of forecasts used in decision-making processes. Finally, these results support a shift in emphasis from pure accuracy to stability-aware model selection and evaluation, in those contexts that are more sensitive to forecast reviews.
Nevertheless, some limitations of the present study should be acknowledged. While the analysis covers a broad range of global models and two state-of-the-art retail forecasting datasets, the generalizability of the findings to other domains or data types (e.g., macroeconomic indicators, financial series, or healthcare time series) remains an open question. Additionally, the study focuses solely on full retraining schemes and does not explore incremental or online learning strategies, which may offer a different trade-off between accuracy and stability. Future research could extend this work by incorporating such adaptive methods, assessing alternative ensembling techniques (e.g., weighted or stacking ensembles), and introducing business-oriented stability metrics that link forecast instability directly to operational cost or risk. Further theoretical development of stability measures, particularly for probabilistic forecasts, may also contribute to a deeper understanding of how models behave across time and under varying data regimes.
In conclusion, this work provides strong empirical evidence that frequent retraining is not necessary for maintaining stability in global forecasting models, and that ensembling, particularly when built on diverse model pools, is a highly effective strategy for achieving both stable and accurate forecasts. These findings contribute to a growing body of research promoting sustainable and robust forecasting practices in modern data-intensive environments.
Supplementary material
In this section, we provide tables and figures related to the empirical results of the M5 and VN1 datasets. Table 4 shows the composition of ensembles across the different datasets.
Table 4: Composition of ensembles for the M5 and VN1 datasets.
Dataset Ensemble Models Included
M5 ENS2A XGBoost, LGBM
ENS3A XGBoost, LGBM, LR
ENS4A XGBoost, LGBM, LR, NBEATSx
ENS5A XGBoost, LGBM, LR, NBEATSx, MLP
VN1 ENS2A MLP, NBEATSx
ENS3A MLP, NBEATSx, NHITS
ENS4A MLP, NBEATSx, NHITS, RF
ENS5A MLP, NBEATSx, NHITS, RF, XGBoost
The Tables 5, 6, 8, and 7 show the forecast instability of the different models and ensembles along the examined retrain scenarios for the M5 daily dataset.
Method 7 14 21 30 60 90 120 150 180 364
LR 0.188 0.188 0.188 0.189 0.189 0.190 0.190 0.192 0.193 0.196
XGBoost 0.075 0.074 0.073 0.073 0.072 0.072 0.072 0.072 0.073 0.073
LGBM 0.054 0.053 0.053 0.053 0.052 0.052 0.052 0.052 0.052 0.052
CatBoost 0.118 0.043 0.059 0.037 0.029 0.020 0.018 0.018 0.018 0.019
MLP 0.500 0.511 0.506 0.520 0.513 0.524 0.510 0.525 0.475 0.553
TCN 0.499 0.209 0.174 0.086 0.055 0.043 0.031 0.017 0.017 0.003
NBEATSx 0.547 0.541 0.527 0.561 0.528 0.536 0.523 0.536 0.515 0.412
NHITS 0.566 0.548 0.558 0.534 0.565 0.537 0.533 0.520 0.512 0.564
Table 5: M5 sMAPC values for each method and retrain scenario combination.
Method 7 14 21 30 60 90 120 150 180 364
Ens2A 0.059 0.058 0.058 0.058 0.058 0.058 0.058 0.058 0.058 0.058
Ens3A 0.089 0.089 0.089 0.089 0.089 0.089 0.090 0.090 0.091 0.092
Ens4A 0.095 0.094 0.094 0.097 0.095 0.095 0.095 0.095 0.096 0.097
Ens5A 0.099 0.098 0.098 0.101 0.099 0.099 0.099 0.099 0.100 0.101
Table 6: M5 sMAPC values for each ensemble and retrain scenario combination.
Method 7 14 21 30 60 90 120 150 180 364
LR 0.118 0.118 0.118 0.119 0.119 0.119 0.119 0.119 0.119 0.119
XGBoost 0.115 0.115 0.115 0.116 0.115 0.115 0.115 0.116 0.115 0.114
LGBM 0.123 0.122 0.123 0.124 0.123 0.123 0.123 0.124 0.123 0.122
CatBoost 0.127 0.144 0.123 0.144 0.144 0.144 0.144 0.142 0.141 0.139
MLP 0.109 0.109 0.110 0.110 0.111 0.111 0.112 0.111 0.110 0.110
TCN 0.114 0.114 0.114 0.113 0.113 0.112 0.113 0.112 0.110 0.108
NBEATSx 0.110 0.109 0.110 0.115 0.110 0.112 0.111 0.111 0.129 0.169
NHITS 0.112 0.111 0.122 0.112 0.114 0.114 0.115 0.135 0.114 0.115
Table 7: M5 MQC values for each method and retrain scenario combination.
Method 7 14 21 30 60 90 120 150 180 364
Ens2A 0.117 0.117 0.117 0.118 0.118 0.118 0.118 0.119 0.118 0.117
Ens3A 0.117 0.117 0.116 0.118 0.118 0.118 0.117 0.118 0.117 0.117
Ens4A 0.113 0.113 0.113 0.114 0.114 0.113 0.113 0.114 0.113 0.113
Ens5A 0.111 0.111 0.111 0.112 0.111 0.111 0.111 0.111 0.110 0.110
Table 8: M5 MQC values for each ensemble and retrain scenario combination.
The Tables 9, 10, 12, and 11 show the forecast instability of the different models and ensembles along the examined retrain scenarios for the VN1 weekly dataset.
Method 1 2 3 4 6 8 10 13 26 52
LR 0.085 0.084 0.085 0.082 0.083 0.086 0.086 0.085 0.084 0.082
RF 0.092 0.090 0.089 0.089 0.088 0.088 0.089 0.089 0.089 0.090
XGBoost 0.146 0.129 0.126 0.120 0.118 0.115 0.114 0.119 0.117 0.113
LGBM 0.121 0.110 0.111 0.106 0.104 0.105 0.105 0.104 0.104 0.105
CatBoost 0.320 0.198 0.186 0.152 0.134 0.139 0.115 0.117 0.100 0.095
MLP 0.458 0.469 0.468 0.475 0.443 0.483 0.438 0.490 0.469 0.452
LSTM 0.558 0.301 0.203 0.126 0.101 0.044 0.055 0.048 0.032 0.013
TCN 0.600 0.260 0.226 0.177 0.120 0.091 0.052 0.056 0.020 0.013
NBEATSx 0.596 0.594 0.596 0.577 0.577 0.581 0.573 0.580 0.571 0.550
NHITS 0.605 0.589 0.583 0.572 0.533 0.541 0.552 0.528 0.573 0.580
Table 9: VN1 sMAPC values for each method and retrain scenario combination.
Method 7 14 21 30 60 90 120 150 180 364
Ens2A 0.503 0.515 0.505 0.495 0.499 0.504 0.487 0.485 0.495 0.509
Ens3A 0.514 0.526 0.526 0.509 0.509 0.514 0.510 0.512 0.511 0.512
Ens4A 0.167 0.170 0.162 0.153 0.145 0.137 0.148 0.134 0.135 0.145
Ens5A 0.164 0.162 0.154 0.145 0.138 0.131 0.138 0.129 0.129 0.135
Table 10: VN1 sMAPC values for each ensemble and retrain scenario combination.
Method 1 2 3 4 6 8 10 13 26 52
LR 1.064 1.067 1.063 1.061 1.070 1.079 1.103 1.097 1.127 1.163
RF 1.322 1.329 1.328 1.341 1.353 1.366 1.403 1.392 1.439 1.504
XGBoost 1.101 1.072 1.080 1.079 1.083 1.097 1.084 1.135 1.175 1.254
LGBM 1.282 1.179 1.235 1.161 1.165 1.173 1.199 1.277 1.266 1.379
CatBoost 1.435 1.281 1.244 1.159 1.150 1.166 1.151 1.225 1.207 1.246
MLP 0.844 0.876 0.887 0.883 0.889 0.891 0.876 0.933 0.903 0.969
LSTM 1.020 1.015 1.008 1.004 1.005 0.993 0.986 0.987 0.972 0.930
TCN 1.020 1.015 1.008 1.004 1.005 0.993 0.986 0.987 0.972 0.930
NBEATSx 1.532 1.519 1.467 1.253 1.167 1.210 1.196 1.069 1.084 1.147
NHITS 1.592 1.505 1.412 1.254 1.212 0.998 1.174 1.028 1.082 1.145
Table 11: VN1 MQC values for each method and retrain scenario combination.
Method 7 14 21 30 60 90 120 150 180 364
Ens2A 1.079 1.114 1.062 0.974 0.949 0.982 0.962 0.921 0.944 1.022
Ens3A 1.200 1.192 1.148 1.024 0.990 0.973 0.995 1.005 0.970 1.034
Ens4A 1.180 1.166 1.122 1.036 1.021 1.001 1.028 0.988 1.017 1.080
Ens5A 1.153 1.137 1.107 1.033 1.022 1.000 1.031 1.006 1.029 1.091
Table 12: VN1 MQC values for each ensemble and retrain scenario combination.
Figures 9, and 10 show the results of the Friedman-Nemenyi test in the context of probabilistic instability for the M5 and VN1 datasets.
Figure 9: M5 Friedman-Nemenyi test results based on MQC.
Figure 10: VN1 Friedman-Nemenyi test results based on MQC.
Figures 11, 12, 13, and 14 show the results of the Friedman-Nemenyi test in the context of both point and probabilistic instability for the different ensemble models across retrain scenarios and for each dataset.
Figure 11: M5 Friedman-Nemenyi test results on ensembles based on sMAPC.
Figure 12: M5 Friedman-Nemenyi test results on ensembles based on MQC.
Figure 13: VN1 Friedman-Nemenyi test results on ensembles based on sMAPC.
Figure 14: VN1 Friedman-Nemenyi test results on ensembles based on MQC.

### v2

@@
-On the stability of global forecasting models
+
 Abstract
-Forecast stability, that is the consistency of predictions over time, is essential in business settings where sudden shifts in forecasts can disrupt planning and erode trust in predictive systems.
-Despite its importance, stability is often overlooked in favor of accuracy, particularly in global forecasting models.
-In this study, we evaluate the stability of point and probabilistic forecasts across different retraining frequencies and ensemble strategies using two large retail datasets (M5 and VN1).
-To do this, we introduce a new metric for probabilistic stability (MQC) and analyze ten different global models and four ensemble configurations.
-The results show that less frequent retraining not only preserves but often improves forecast stability, while ensembles, especially those combining diverse pool of models, further enhance consistency without sacrificing accuracy.
-These findings challenge the need for continuous retraining and highlight ensemble diversity as a key factor in reducing forecast stability.
-The study promotes a shift toward stability-aware forecasting practices, offering practical guidelines for building more robust and sustainable prediction systems.
+Forecast stability, that is, the consistency of predictions over time, is essential in business settings where sudden shifts in forecasts can disrupt planning and erode trust in predictive systems.
+Despite its importance, stability is often overlooked in favor of accuracy.
+In this study, we evaluate the stability of point and probabilistic forecasts across several retraining scenarios using two large retail demand datasets (M5 and VN1) and ten different global forecasting models.
+To analyze stability in the probabilistic setting, we propose a new model-agnostic, distribution-free, and scale-free metric that measures probabilistic instability: the Scaled Multi-Quantile Change (SMQC).
+Furthermore, we also evaluate the effects of retraining on various ensemble configurations based on forecast pooling.
+The results show that, compared to continuous retraining, less frequent retraining not only preserves but often improves forecast stability, challenging the need for continuous retraining.
+The study promotes a shift toward stability-aware forecasting practices, proposing a new tool to effectively evaluate forecast stability in probabilistic settings, and offering practical guidelines for building more robust prediction systems.
 keywords:
@@
 In many operational contexts, forecasts are produced regularly and serve as the foundation for critical business decisions, from inventory planning to resource allocation.
-In such settings, forecast stability becomes a key requirement; not only should forecasts be accurate, but they should also remain reasonably consistent over time.
+In such settings, forecast stability becomes a key requirement; not only should forecasts be accurate, but they should also remain reasonably stable over time.
 Indeed, forecast stability refers to the consistency of predictions produced by a forecasting model as new data becomes available.
@@
 Both types play a distinct role: vertical stability helps avoid costly forecast revisions that can disrupt planning cycles, while horizontal stability prevents erratic behavior across time steps that could lead to inefficient operations and amplification of demand fluctuations, such as the bullwhip effect in supply chains (Lee et al., 1997).
-In particular, vertical stability may also be seen as temporal robustness, in the sense that a forecasting model is robust to updates (or retraining) as new observations are available, a property that is critical in practice for avoiding frequent changes in business decisions.
+In particular, vertical stability may also be seen as temporal robustness, in the sense that a forecasting model is stable to updates (or retraining) as new observations are available, a property that is critical in practice for avoiding frequent changes in business decisions.
 Forecast instability, indeed, can have serious consequences.
@@
 Conversely, overly stable forecasts might sacrifice accuracy by ignoring valuable new information.
-Ensembling techniques, which integrate multiple forecasts from different models, have been proposed as a mechanism to concurrently reduce bias and variance while enhancing stability (Wang et al., 2023).
+Ensembling techniques, which integrate multiple forecasts from different models, have been proposed as a mechanism to concurrently reduce bias and variance (Wang et al., 2023).
 By aggregating diverse predictions, ensembling methods can alleviate the impact of individual model volatility and potentially achieve a more favorable equilibrium between accuracy and stability.
-Even if the trade-off between forecasting accuracy and stability presents a significant challenge for organizations seeking to balance predictive performance with operational consistency, recent research suggests that stability and accuracy are not necessarily in conflict, and both can be achieved at the same time ((Godahewa et al., 2025), (Van Belle et al., 2023)).
+Even if the trade-off between forecasting accuracy and stability presents a significant challenge for organizations seeking to balance predictive performance with operational stability, recent research suggests that high levels of stability and accuracy can be achieved at the same time ((Godahewa et al., 2025), (Van Belle et al., 2023)).
 Moreover, Zanotti (2025b) demonstrated that the retraining frequency does not harm the forecast accuracy of global models, meaning that models estimated using this forecasting approach retain the same level of performance even when updated less frequently.
-This has a strong potential implication on the forecast stability: if we can reduce the retraining frequency of global models, then we should obtain more stable predictions, because stability should be a non-increasing function of the retraining frequency.
+This has a strong potential implication on the forecast stability: if we can reduce the retraining frequency of global models, then we should obtain more stable predictions, because retraining leads to different model artifacts potentially producing stability disruptions.
 1.1 Research Question
-We aim to address the question ”Are global forecasting models stable?”.
+We aim to address the question ”Do global models for retail demand forecasting produce stable forecasts?”.
 Purposely, we study whether the global modeling approach produces stable forecasts, and we try to understand the effects of retraining on the forecast stability, that is whether avoiding re-estimation for every new observation damages the stability of global models.
@@
   • 2.
-    We analyze the relationship between retraining frequency and forecast stability, comparing different scenarios to quantify the impact of frequent retraining in terms of the stability of forecasting.
+    We analyze the relationship between the retraining period and forecast stability, comparing different scenarios to quantify the impact of frequent retraining in terms of the stability of forecasting.
   • 3.
-    We suggest a new metric to evaluate the instability of probabilistic forecasts.
+    We suggest new metrics to evaluate the instability of probabilistic forecasts.
   • 4.
@@
 Today, most time series forecasting studies include at least a benchmark comparison involving global models (GMs), underscoring their growing importance in the field.
+Following Zanotti (2025b), we define $\mathcal{Y}$ as the set of all available time series in a dataset, where each $Y_{i}\in\mathcal{Y}$ represents an individual time series.
+Let $\mathcal{F}$ denote the set of possible predictive functions, with each $F\in\mathcal{F}$ corresponding to a specific forecasting model.
+Without loss of generality, we assume that all necessary information for prediction is contained within $\mathcal{Y}$.
+Under the local modeling paradigm, forecasts for a horizon $h$ are generated by training a separate model for each individual time series.
+This implies that each series $Y_{i}$ is associated with its own model, characterized by its own set of parameter values.
+$Y_{i}^{h}=F(Y_{i},\theta_{i})\mbox{.}$ (1)
+In contrast, under the global modeling framework, forecasts for each individual time series are generated using a single model trained on the entire dataset.
+This approach leverages cross-series information, allowing the model to learn shared patterns and structures across all time series in $\mathcal{Y}$.
+$Y_{i}^{h}=F(\mathcal{Y},\Theta)\mbox{.}$ (2)
+It is important to note that in the cross-learning methodology, the model parameters $\Theta$ are not specific to individual time series but are shared across all series in the dataset.
+This parameter sharing is a key characteristic of global models, enabling them to generalize patterns across series and potentially reducing the forecast instability.
 Semenoglou et al.
@@
 Recently, research has begun focusing on improving GMs’ ability to capture local patterns (Sen et al., 2019) and enhance their interpretability (Rajapaksha et al., 2023).
-From a forecasting stability perspective, the literature is very poor.
+From a forecasting stability perspective, the literature is impoverished.
 Godahewa et al.
@@
 Van Belle et al.
-(2023), instead, extended an existing deep learning architecture (NBEATS) to optimize forecasts from both a traditional forecast accuracy perspective as well as a forecast stability perspective, directly including an instability component into the loss function of the model.
+(2023), instead, extended an existing deep learning architecture (NBEATS) to optimize forecasts from both a traditional forecast accuracy perspective, as well as a forecast stability perspective, directly including an instability component into the loss function of the model.
 There is also active research on forecast stability within judgmental forecasting ground (Fildes & Goodwin, 2021).
@@
 Nevertheless, in many forecasting applications, such as supply chain management, it is crucial to generate and evaluate predictions probabilistically, whether through prediction intervals, quantiles, or full predictive distributions (Fildes et al., 2022).
-Among the methods developed for uncertainty quantification, Vovk et al.
-(2005) introduced Conformal Inference, a model-agnostic framework that offers valid uncertainty estimates and can also be applied in time series forecasting settings (Stankeviciute et al., 2021).
-In the context of model retraining and updating strategies, the most comprehensive work in time series forecasting is by Zanotti (2025b), who showed the positive effects of reducing the retraining frequency of global models on both accuracy and forecasting costs.
-Spiliotis & Petropoulos (2024) extensively investigated the impact of various retraining scenarios and parameter update methods on model performance.
-However, their focus was limited to the exponential smoothing family, following the traditional local modeling approach.
+The seminal work of Van Belle et al.
+(2024) is the only one addressing the problem of probabilistic stability evaluation, but it only applies to a specific forecasting model (NBEATS), which has been modified to output a full predictive Gaussian distribution.
+Therefore, the stability evaluation of probabilistic forecasts is an open question, especially in a model-agnotic way.
+Lukily, among the methods developed for uncertainty quantification, Vovk et al.
+(2005) introduced Conformal Inference, which is a model-agnostic framework that offers valid uncertainty estimates and can also be applied in time series forecasting settings (Stankeviciute et al., 2021).
+In the context of model retraining and updating strategies, the most comprehensive work in time series forecasting is by Zanotti (2025b), who showed the positive effects of reducing retraining of global models on both accuracy and forecasting costs.
+The seminal work of Spiliotis & Petropoulos (2024) extensively investigated the impact of various retraining scenarios and parameter update methods on model performance, focusing on the exponential smoothing family of models, following the traditional local modeling approach.
 (Huber & Stuckenschmidt, 2020) touched on retraining within the retail demand domain, but the study was restricted to a small set of models, limited retraining strategies, and a proprietary daily dataset.
@@
 The M5 competition, part of the well-known M-competitions series organized by Spyros Makridakis and colleagues, aimed to benchmark forecasting methods in the context of retail demand (Makridakis et al., 2022b).
-The M5 dataset (Howard & Makridakis, 2020) includes 3,049 daily time series representing unit sales of Walmart products across three categories (Food, Hobbies, and Household) sold in ten stores located in California, Texas, and Wisconsin.
+The M5 dataset (Howard & Makridakis, 2020) includes 30,490 daily time series representing unit sales of Walmart products across three categories (Food, Hobbies, and Household) sold in ten stores located in California, Texas, and Wisconsin.
 The data span from 2011 to 2016 and are characterized by high intermittency and a hierarchical structure, enabling forecasts at multiple aggregation levels (e.g., SKU, product category, store, and state).
@@
 Together, these two datasets offer the most recent and comprehensive time series collections related to retail demand, enabling good generalizability of our findings.
-Table 1: The M5 and the VN1 datasets used in the experiments.
-Dataset Frequency N. Series Min Obs per Series
-M5 Daily (7) 28.298 730
-VN1 Weekly (52) 15.053 157
-In both cases, our analysis focused on the most disaggregated level of the data (i.e., SKUs), where the potential advantages of reducing retraining frequency are likely to be most pronounced.
+In both cases, our analysis focused on the most disaggregated level of the data (i.e., SKUs), where the potential advantages of retraining less frequently are likely to be most pronounced.
 To ensure consistency with the evaluation setup described in Section 3.3, we applied a filtering criterion: for the daily data (M5), only time series with at least two years of observations (730 data points) were retained; for the weekly data (VN1), we considered only those series with at least three years of data (157 observations).
@@
 In this section, we present an overview of the global models used in our experiments.
-Following Zanotti (2025b), let us define $\mathcal{Y}$ as the set of all available time series in a dataset, where each $Y_{i}\in\mathcal{Y}$ represents an individual time series.
-Let $\mathcal{F}$ denote the set of possible predictive functions, with each $F\in\mathcal{F}$ corresponding to a specific forecasting model.
-Without loss of generality, we assume that all necessary information for prediction is contained within $\mathcal{Y}$.
-Under the local modeling paradigm, forecasts for a horizon $h$ are generated by training a separate model for each individual time series.
-This implies that each series $Y_{i}$ is associated with its own model, characterized by its own set of parameter values.
-$Y_{i}^{h}=F(Y_{i},\theta_{i})\mbox{.}$ (1)
-In contrast, under the global modeling framework, forecasts for each individual time series are generated using a single model trained on the entire dataset.
-This approach leverages cross-series information, allowing the model to learn shared patterns and structures across all time series in $\mathcal{Y}$.
-$Y_{i}^{h}=F(\mathcal{Y},\Theta)\mbox{.}$ (2)
-It is important to note that in the cross-learning methodology, the model parameters $\Theta$ are not specific to individual time series but are shared across all series in the dataset.
-This parameter sharing is a key characteristic of global models, enabling them to generalize patterns across series and potentially reducing the forecast instability.
-In our study, we focused exclusively on analyzing the stability of global methods, as our primary goal was to assess whether this modeling approach, unlike the traditional local one, can maintain performance with less frequent retraining.
+In our study, we focused exclusively on analyzing the stability of global methods, as our primary goal was to assess whether this modeling approach, unlike the traditional local one, can maintain performance with less frequent retraining 111 The analysis of the stability of local models was out of the scope of our study.
+.
 Cross-learning has become the standard in many industries dealing with large-scale time series data, particularly in retail demand forecasting, where regular forecasts for thousands of SKUs are required (Januschowski et al., 2020).
@@
 The selected models are well-established in the time series forecasting literature and represent a range of methodological paradigms, allowing for a broad and informative comparison.
-As in Zanotti (2025b), we used five machine learning models and five deep learning models.
 Machine learning models have proven effective in forecasting tasks due to their ability to capture complex non-linear relationships in data.
@@
 We constructed time series features such as lags, rolling means, expanding means, and calendar features (e.g., year, month, week, day of the week).
-We also included static metadata such as store, product, category, and location identifiers based on the dataset’s frequency.
+We also included static metadata such as store, product, category, and location identifiers.
 For the M5 dataset, additional external features like special events were also incorporated.
@@
 Nevertheless, training deep models is often more complex due to the higher number of hyperparameters and their sensitivity to configuration choices, which can significantly affect performance (Smyl, 2020).
-In our implementation, we followed competition-proven practices by providing only static, calendar, and external covariates as inputs, while using top solutions’ guidelines for setting the model hyperparameters.
+In our implementation, we followed competition-proven practices by providing only static, calendar, and external covariates, together with historical values, as inputs.
+Top solutions’ guidelines were adopted for setting the model hyperparameters when available, and otherwise, the recommended defaults from the respective libraries were used.
 We compared five deep learning models for time series forecasting.
@@
 Neural Basis Expansion Analysis for Time Series (NBEATS) and Neural Hierarchical Interpolation for Time Series (NHITS) are state-of-the-art models for time series forecasting, with NBEATS offering interpretable trend and seasonality components (Oreshkin et al., 2020), and NHITS improving upon NBEATS with hierarchical interpolation mechanisms (Challu et al., 2022).
-Moreover, we also estimated four different ensemble learning models based on Zanotti (2025a).
-Ensemble learning combines the predictions of multiple base models to improve forecast accuracy and robustness, particularly when individual models capture complementary patterns in the data.
+Ensemble learning combines the predictions of multiple base models to improve forecast accuracy, particularly when individual models capture complementary patterns in the data.
 In the context of global forecasting, ensembles are especially valuable for mitigating the risk of model instability and overfitting, which can occur when relying on a single method (Wang et al., 2023).
-In our study, we adopted a simple averaging strategy, which is widely used for its ease of implementation and effectiveness in improving forecast robustness (Claeskens et al., 2016).
-Specifically, we constructed ensembles by averaging the forecasts of the top two, top three, top four, and top five most accurate individual models, as determined on out-of-sample performance.
-The ensemble configurations were designed to reflect practical setups and were validated using the same rolling origin evaluation framework described in Section 3.3.
-By incorporating ensemble learning into our analysis, we aimed to assess whether combining global models could further enhance forecasting stability under different retraining scenarios.
-Moreover, by evaluating multiple ensemble sizes, we aimed to understand how the inclusion of additional models affects the stability under different circumstances.
+Following Zanotti (2025a), we estimated four different ensemble models, adopting a simple averaging strategy, which is widely used for its ease of implementation and effectiveness in improving forecast robustness (Claeskens et al., 2016).
+Indeed, simple averaging can be used for both point and probabilistic forecasting, where equally weighting quantiles has been shown to yield robust results Busetti (2017).
+To evaluate the potential benefits of ensembling, we simulated a range of ensemble models to assess how forecast stability across the full test window would have differed if combinations had been employed in place of individual base models.
+Specifically, we constructed simple mean ensembles by aggregating the forecasts of subsets of global models, selecting only those base learners that achieved the highest individual accuracy.
+This design leverages the strengths of top-performing models while aiming to mitigate their individual weaknesses through averaging.
+Ensembles of increasing size were considered, combining the top 2, 3, 4, and 5 models according to each out-of-sample performance.
+Such selective pooling is essentially a variant of model pooling in the sense of Kourentzes et al.
+(2019), where only a subset of candidate models contributes to the final forecast.
+Nevertheless, to maintain a balance between potential improvements and the diminishing returns or additional complexity typically associated with larger ensembles (Wang et al., 2023), we limited the maximum ensemble size to five models (out of the eight or ten available, depending on the dataset).
+The ensemble models were finally validated using the same rolling origin evaluation framework described in Section 3.3.
+By incorporating ensemble learning into our analysis, we aimed to study whether combining global models could further enhance forecasting stability under different retraining scenarios.
 All global models were implemented in Python using Nixtla’s framework (Nixtla, 2022).
@@
 The model is then retrained on the updated training data, using either a fixed-length or expanding window.
-Stability metrics (see Section 3.4) are averaged over all iterations to provide a robust estimate of forecasting accuracy.
+Evaluation metrics (see Section 3.4) are averaged over all iterations to provide a robust estimate of forecasting accuracy.
 Compared to fixed origin evaluations, rolling origin offers a more nuanced view of a model’s robustness by exposing it to a range of temporal conditions (including seasonal patterns, trends, and potential anomalies) that may not be captured in a single evaluation window (Bergmeir & Benítez, 2012).
@@
 This approach is widely used in applied settings (Petropoulos & et al., 2022), and it is the one we adopted in our study, especially given the relatively short length of the weekly VN1 series.
-Following Zanotti (2025b), Table 2 outlines the key parameters used in our experiments, including a test set spanning at least one full year to mitigate seasonal bias, horizons aligned with typical business use cases, and a step size of one to maximize evaluation frequency.
-Table 2: The retraining scenarios, the test window, and the horizon for the M5 and the VN1 datasets.
-Dataset Frequency Retraining Scenarios (r) Test Window (T) Horizon (h)
-M5 Daily (7) 7, 14, 21, 30, 60, 90, 120, 150, 180, 364 364 28
-VN1 Weekly (52) 1, 2, 3, 4, 6, 8, 10, 13, 26, 52 52 13
-Furthermore, to investigate the impact of retraining frequency on forecasting stability, we examined various retraining scenarios, or retraining windows, as defined in Zanotti (2025b).
-Each window $r$ indicates the number of new observations after which the model is retrained.
+The Table 1 outlines the key parameters used in our experiments, including a test set spanning at least one full year, to mitigate seasonal bias, horizons aligned with typical business use cases, and a step size of one to maximize evaluation frequency.
+A step size of one is also relevant in the context of stability evaluation, since multiple forecasts for the same target are required to compute stability metrics.
+Table 1: Retraining set, train and test size, frequency, forecast horizon, and step size for each dataset. The setup is based on the respective time series frequency.
+M5 VN1
+Frequency ($f$) daily (7) weekly (52)
+Train size ($n$) $\geq 730$ $\geq 157$
+Test size ($T$) 364 52
+Horizon ($h$) 28 13
+Step size ($p$) 1 1
+Retrain set ($r$) {7, 14, 21, 30, 60, 90, 120, 150, 180, 364} {1, 2, 3, 4, 6, 8, 10, 13, 26, 52}
+Furthermore, to investigate the impact of retraining on forecasting stability, we examined various retraining scenarios, or retraining windows, as defined in Zanotti (2025b).
+Each scenario $r$ indicates the number of new observations after which the model is retrained.
 These retraining scenarios are tailored to the data frequency, daily for M5 and weekly for VN1, since frequency dictates both the forecast horizon and business review cycles.
-Table 2 summarizes the selected retraining windows, test periods, and forecast horizons.
 For instance, in the daily case, $r=7$ reflects weekly updates.
@@
 $\text{sMAPC}=\frac{200}{h-1}\sum_{t=n+1}^{n+h-1}\frac{\lvert\hat{y}_{t,n}-\hat{y}_{t,n-1}\rvert}{\lvert\hat{y}_{t,n}\rvert-\lvert\hat{y}_{t,n-1}\rvert}\mbox{.}$ (3)
-Here, $\hat{y}_{t,n}$ and $\hat{y}_{t,n-1}$ are the forecasts generated for period $t$ with origins $n$ and $n-1$ respectively.
+Here, $\hat{y}_{t,n}$ and $\hat{y}_{t,n-1}$ are the forecasts generated for period $t$ with origins $n$ and n1n-1 respectively.
 The instability across different pairs of consecutive forecasting origins can be obtained by a simple average of the sMAPC values.
@@
 Since most machine learning and deep learning models do not natively produce full predictive distributions, we employed Conformal Inference to construct prediction intervals around point forecasts.
-Conformal Inference is a flexible and robust framework that provides distribution-free, model-agnostic prediction intervals with guaranteed finite-sample coverage under mild assumptions (Vovk et al., 2005).
-Although initially developed for i.i.d.
-data, recent extensions have adapted it for time series applications by relaxing the exchangeability assumption (Stankeviciute et al., 2021).
-Its advantages (validity guarantees, minimal assumptions, computational simplicity, and effectiveness with limited data) make it especially suitable for our global forecasting setup, where different models are compared across heterogeneous datasets and varying levels of retraining frequency.
-Because no metric exists to evaluate the vertical stability (or instability) of the resulting probabilistic forecasts, we proposed a new measure, the Multi-Quantile Change (MQC, or Multi-Quantile Loss Change, MQLC), to comprehensively assess the stability of the probabilistic predictions.
+Its advantages (validity guarantees, minimal assumptions, computational simplicity, and effectiveness with limited data) make it especially suitable for our global forecasting setup, where different models are compared across heterogeneous datasets and varying retraining scenarios 222 One may prefer, instead, to train models using quantile losses to directly yield quantile forecasts, or to use inherently probabilistic models.
+These are also valid ways of modeling uncertainty, as demonstrated by the M5 Uncertainty competition (Makridakis et al., 2022c; Lainder & Wolfinger, 2022).
+However, comparing different probabilistic modeling frameworks is out of the scope of our analysis, and it may be an area of future research.
+.
+Because no metric exists to evaluate the vertical stability (or instability) of the resulting probabilistic quantile forecasts, we proposed a new measure, the Multi-Quantile Change (MQC), to comprehensively assess the stability of the probabilistic predictions.
 We defined the Quantile Change and the Multi-Quantile Change as:
-$\text{QC}=\frac{1}{h-1}\sum_{t=n+1}^{n+h-1}\left(q\cdot(\hat{y}_{t,n-1}-\hat{y}_{t,n})\cdot\mathbb{I}_{\hat{y}_{t,n-1}\geq\hat{y}_{t,n}}+(1-q)\cdot(\hat{y}_{t,n}-\hat{y}_{t,n-1})\cdot\mathbb{I}_{\hat{y}_{t,n-1}<\hat{y}_{t,n}}\right)\mbox{,}$ (4)
-$\text{MQC}=\frac{1}{\mathcal{Q}}\sum_{q\in\mathcal{Q}}\text{QC}(q)\mbox{.}$ (5)
-The Quantile Change (QC) metric is a measure of the change in forecasted quantiles between two consecutive forecast origins.
-Given two forecast origins, $n$ and $n-1$, QC quantifies the average adjustment in the predicted quantiles across the forecast horizon $h$.
-The formula mirrors the Quantile Loss (Pinball Loss) but replaces the actual observations $y_{t}$, with previous forecasts $\hat{y}_{t,n-1}$, thereby shifting the focus from forecast accuracy to forecast stability.
-The intuition is straightforward: if the forecasts are stable across time (i.e., $\hat{y}_{t,n}\approx\hat{y}_{t,n-1}$), then QC will be small.
-Large QC values indicate that the model updates its distribution significantly with each new observation, suggesting a lack of temporal robustness in its probabilistic output.
-By aggregating this across a range of quantiles $\mathcal{Q}$, the Multi-Quantile Change (MQC) captures overall instability across the entire forecast distribution, providing a comprehensive metric for probabilistic stability.
-One drawback of the MQC is that it is based solely on a finite set of quantiles rather than the full predictive distribution.
-As a result, it captures changes only at selected points in the distribution (e.g., median, tails), potentially missing more subtle shifts in shape, skewness, or variance that occur between those quantiles.
+$\text{QC}(\alpha)=\frac{1}{h-1}\sum_{t=n+1}^{n+h-1}\Big{(}\alpha\cdot(\hat{q}_{t,n-1}^{(\alpha)}-\hat{q}_{t,n}^{(\alpha)})\cdot\mathbb{I}_{\hat{q}_{t,n}^{(\alpha)}\leq\hat{q}_{t,n-1}^{(\alpha)}}+(1-\alpha)\cdot(\hat{q}_{t,n}^{(\alpha)}-\hat{q}_{t,n-1}^{(\alpha)})\cdot\mathbb{I}_{\hat{q}_{t,n}^{(\alpha)}>\hat{q}_{t,n-1}^{(\alpha)}}\Big{)}\mbox{,}$ (4)
+$\text{MQC}=\frac{1}{|\mathcal{Q}|}\sum_{\alpha\in\mathcal{Q}}\text{QC}(\alpha)\mbox{,}$ (5)
+where $\mathcal{Q}$ is the set of possible quantile levels $\alpha$.
+The Quantile Change (QC) metric is a measure of the change in forecasted quantiles, $\hat{q}_{t}^{(\alpha)}$, between two consecutive forecast origins.
+Given two forecast origins, $n$ and n1n-1, QC quantifies the average adjustment in the predicted quantiles across the forecast horizon $h$.
+The formula mirrors the Quantile Loss (Pinball Loss) but replaces the actual observations $y_{t}$, with previous quantile forecasts $\hat{q}_{t,n-1}^{(\alpha)}$, thereby shifting the focus from forecast accuracy to forecast stability.
+The intuition is straightforward: if the forecasts are stable across time (i.e., $\hat{q}_{t,n}^{(\alpha)}\approx\hat{q}_{t,n-1}^{(\alpha)}$), then QC will be small.
+Large QC values indicate that the model updates its distribution significantly with each new observation, suggesting a lack of stability in its probabilistic output.
+By aggregating this across a set of quantile levels $\mathcal{Q}$, the Multi-Quantile Change (MQC) captures overall instability, providing a comprehensive metric for probabilistic stability.
+One drawback of the MQC is that it is a scale-dependent measure of instability. To solve this issue, we also propose a scaled version of the MQC, the Scaled Multi-Quantile Change (SMQC), defined as
+$\text{SQC}(\alpha)=\frac{\frac{1}{h-1}\sum_{t=n+1}^{n+h-1}\Big{(}\alpha\cdot(\hat{q}_{t,n-1}^{(\alpha)}-\hat{q}_{t,n}^{(\alpha)})\cdot\mathbb{I}_{\hat{q}_{t,n}^{(\alpha)}\leq\hat{q}_{t,n-1}^{(\alpha)}}+(1-\alpha)\cdot(\hat{q}_{t,n}^{(\alpha)}-\hat{q}_{t,n-1}^{(\alpha)})\cdot\mathbb{I}_{\hat{q}_{t,n}^{(\alpha)}>\hat{q}_{t,n-1}^{(\alpha)}}\Big{)}}{\frac{1}{n-s}\sum_{t=s+1}^{n}\left|y_{t}-y_{t-s}\right|}\mbox{,}$ (6)
+$\text{SMQC}=\frac{1}{|\mathcal{Q}|}\sum_{\alpha\in\mathcal{Q}}\text{SQC}(\alpha)\mbox{,}$ (7)
+where we adopted the in-sample one-step-ahead seasonal naive forecast as a scaling factor, to enable the effective application of this metric to forecasting models by ensuring that the results are not biased by the scale of individual time series.
+To further contextualize the proposed metric, it is important to compare it with existing distributional distance metrics such as the Kullback–Leibler ($KL$) divergence and the Wasserstein distance ($W_{2}$).
+KL divergence measures the relative entropy between two probability distributions and is particularly sensitive to differences in the tails (Kullback & Leibler, 1951).
+However, it requires access to full predictive densities, which can limit its use in forecasting applications where outputs are often expressed only as quantiles.
+Wasserstein distance ($W_{2}$), by contrast, quantifies the minimal cost of transforming one distribution into another and effectively captures both location and shape differences, but it again relies on full distributions (Peyré & Cuturi, 2019).
+The MQC metric addresses a complementary objective: it evaluates the stability of probabilistic forecasts directly at the quantile level, without the need of full predictive densities.
+Since it is based solely on a finite set of quantiles rather than the full predictive distribution, MQC does not capture all aspects of the distributional change, potentially missing more subtle shifts in shape, skewness, or variance that occur between those quantiles.
 However, in practical forecasting applications, especially in retail and supply chain domains where decisions are often made based on specific quantile levels (e.g., high quantiles for safety stock), this limitation is mitigated by the fact that the selected quantiles are usually those of greatest operational interest.
-Moreover, through Conformal Inference, it is possible to calculate as many quantiles as desired at no computational cost, making MQC a pragmatic and model-agnostic metric for capturing distributional instability across time.
-We computed 13 quantiles: the median and 6 central prediction intervals (60%, 70%, 80%, 90%, 95% and 99%).
-Central intervals describe the forecast center, while wider intervals capture tail risks, essential for safety stock decisions in retail (Barrow & Kourentzes, 2016).
-To ensure reliable quantile estimates, conformal prediction intervals were computed on a validation set at least twice the forecast horizon, which constrained the number of time series used.
-We normalized each evaluation metric relative to the baseline retraining scenario, defined by the dataset frequency, to enable consistent comparison across models and retraining windows.
+Additionally, given that it is based on the pinball loss, the MQC is inherently asymmetric, reflecting the fact that quantile forecasts are directional.
+For example, an unstable 90% quantile forecast may have stronger implications on inventory or risk management than a similarly unstable 60% quantile forecast.
+This implies that MQC emphasizes directional changes when measuring stability, in contrast to symmetric distances (like $W_{2}$), which treat deviations identically regardless of the quantile’s position.
+Importantly, MQC is both model-agnostic and distribution-free, meaning that it can be applied consistently across different forecasting approaches without requiring strong assumptions about the underlying predictive distributions.
+Moreover, MQC is computationally lightweight compared to the Wasserstein distance or the KL divergence, which makes it suitable for large-scale applications involving thousands of time series.
+We considered a total of 14 different quantiles of probability levels $\alpha\in$ {0.005, 0.025, 0.050, 0.100, 0.150, 0.200, 0.025, 0.750, 0.800, 0.850, 0.900, 0.950, 0.975, 0.995}.
+Therefore, we analyzed the 50% 60%, 70%, 80%, 90%, 95%, and 99% central prediction intervals.
+The 50% and 60% central prediction intervals provide a good description of the center of the forecast distribution, while the 90%, 95%, and 99% give more information about its tails, essential to determine the safety stock levels in retail demand forecasting (Barrow & Kourentzes, 2016).
+These quantiles provide sufficient information about the uncertainty of the forecasts and allow for the effective description of the whole distribution.
+To ensure reliable quantile estimates, conformal prediction intervals were computed on a validation set at least twice the forecast horizon, which constrained the number of time series used 333 We used a validation set four times larger than the forecast horizon for the M5 dataset and twice the forecast horizon for the VN1 dataset.
+This limited the number of time series from 30,490 to 28,298 for the M5 dataset.
+.
+We also normalized each evaluation metric relative to the baseline retraining scenario, defined by the dataset frequency, to enable consistent comparison across models and retraining periods.
 To statistically validate our findings, we applied the Friedman-Nemenyi test for multiple comparisons (Demšar, 2006).
@@
 $\text{RMSSE}=\sqrt{\frac{\frac{1}{h}\sum_{t=n+1}^{n+h}(y_{t}-\hat{y}_{t})^{2}}{\frac{1}{n-s}\sum_{t=s+1}^{n}(y_{t}-y_{t-s})^{2}}}\mbox{.}$ (8)
-which compares the model’s squared error to that of a seasonal naive forecast, and it was the official accuracy metric in the M5 competition (Makridakis et al., 2022b).
+which compares the model’s squared error to that of a seasonal naive forecast, and it was the official accuracy metric in the M5 competition (with $s=1$) (Makridakis et al., 2022b).
 Lower values indicate better performance.
-To assess probabilistic accuracy, we evaluated the quantiles produced through the Conformal Inference framework using the Quantile Loss (QL) and Multi-Quantile Loss (MQL).
-$\text{QL}=\frac{1}{h}\sum_{t=n+1}^{n+h}\left(q\cdot(y_{t}-\hat{y}_{t})\cdot\mathbb{I}_{y_{t}\geq\hat{y}_{t}}+(1-q)\cdot(\hat{y}_{t}-y_{t})\cdot\mathbb{I}_{y_{t}<\hat{y}_{t}}\right)\mbox{,}$ (7)
-$\text{MQL}=\frac{1}{\mathcal{Q}}\sum_{q\in\mathcal{Q}}\text{QL}(q)\mbox{.}$ (8)
-These proper scoring rules assess forecast distribution accuracy (Kolassa, 2016). Moreover, the MQL was the official metric of the M5 Uncertainty competition (Makridakis et al., 2022c).
+To assess probabilistic accuracy, we evaluated the quantiles produced using a scaled version of the Multi-Quantile Loss.
+The Scaled Quantile Loss, also known as (Scaled) Pinball Loss, and the Scaled Multi-Quantile Loss are defined as:
+$\text{SQL}(\alpha)=\frac{\frac{1}{h}\sum_{t=n+1}^{n+h}\Big{(}\alpha\cdot(y_{t}-\hat{q}_{t}^{(\alpha)})\cdot\mathbb{I}_{\hat{q}_{t}^{(\alpha)}\leq y_{t}}+(1-\alpha)\cdot(\hat{q}_{t}^{(\alpha)}-y_{t})\cdot\mathbb{I}_{\hat{q}_{t}^{(\alpha)}>y_{t}}\Big{)}}{\frac{1}{n-s}\sum_{t=s+1}^{n}|y_{t}-y_{t-s}|}\mbox{,}$ (9)
+$\text{SMQL}=\frac{1}{|\mathcal{Q}|}\sum_{\alpha\in\mathcal{Q}}\text{SQL}(\alpha)\mbox{,}$ (10)
+These proper scoring rules assess forecast distribution accuracy (Kolassa, 2016).
+Moreover, the SMQL was the official metric of the M5 Uncertainty competition (weighted and with $s=1$) (Makridakis et al., 2022c).
 For this study, we adopted a cloud computing machine NC6s_v3 hosted on Microsoft Azure, with Linux Ubuntu 24 operating system, 1 Graphical Processing Unit, 6 Computing Processing Units, 112GB of memory.
 4 Results and discussion
-This section discusses the empirical findings of our study, highlighting the interplay between accuracy, probabilistic performance, and stability cost across different retraining strategies and ensemble configurations.
-We extract insights from both the M5 and VN1 datasets to uncover consistent patterns as well as nuances unique to each dataset’s characteristics.
-Table 3: Overall forecasting performance and stability for each method across datasets. RMSSE, MQL, sMAPC, and MQC. Minimum values per column are highlighted in bold.
+This section discusses the empirical findings of our study, highlighting the interplay between accuracy and stability across different retraining strategies.
+Table 2: Overall forecasting performance and stability for each method across datasets based on the baseline retraining scenario, $r=7$ for M5 and $r=1$ for VN1.
+Minimum values per column are highlighted in bold.
 Method M5 VN1
-RMSSE MQL sMAPC MQC RMSSE MQL sMAPC MQC
-LR 0.777 0.267 0.188 0.118 6.549 2.896 0.085 1.064
-RF 1.868 2.590 0.092 1.322
-XGBoost 0.755 0.258 0.075 0.115 1.890 2.469 0.146 1.101
-LGBM 0.771 0.256 0.054 0.123 3.542 2.625 0.121 1.282
-CatBoost 0.947 0.263 0.118 0.127 5.762 2.845 0.320 1.435
-MLP 0.821 0.281 0.500 0.109 1.543 2.492 0.458 0.844
-LSTM 1.913 2.843 0.558 1.020
-TCN 0.865 0.290 0.499 0.114 1.913 2.843 0.600 1.020
-NBEATSx 0.815 0.279 0.547 0.110 1.698 2.626 0.596 1.532
-NHITS 0.828 0.284 0.566 0.112 1.699 2.632 0.605 1.592
-Ens2A 0.757 0.255 0.059 0.117 1.472 2.369 0.503 1.079
-Ens3A 0.757 0.256 0.089 0.117 1.517 2.410 0.514 1.200
-Ens4A 0.758 0.249 0.095 0.113 1.524 2.386 0.167 1.180
-Ens5A 0.763 0.251 0.099 0.111 1.544 2.375 0.164 1.153
+RMSSE SMQL sMAPC SMQC RMSSE SMQL sMAPC SMQC
+LR 0.777 0.267 0.188 0.110 6.549 4.286 0.085 1.560
+RF 1.868 1.220 0.092 0.729
+XGBoost 0.755 0.245 0.075 0.100 1.890 1.503 0.146 0.629
+LGBM 0.771 0.246 0.054 0.118 3.542 2.318 0.121 1.484
+CatBoost 0.947 0.277 0.118 0.156 5.762 4.233 0.320 2.276
+MLP 0.821 0.263 0.500 0.083 1.543 0.806 0.458 0.219
+LSTM 1.913 0.892 0.558 0.242
+TCN 0.865 0.270 0.499 0.083 1.913 0.891 0.600 0.241
+NBEATSx 0.815 0.262 0.547 0.084 1.698 0.991 0.596 0.765
+NHITS 0.828 0.266 0.566 0.087 1.699 0.992 0.605 0.755
+Ens2A 0.757 0.243 0.059 0.108 1.472 0.826 0.503 0.441
+Ens3A 0.757 0.248 0.089 0.107 1.517 0.867 0.514 0.539
+Ens4A 0.758 0.240 0.095 0.099 1.524 0.926 0.167 0.545
+Ens5A 0.763 0.241 0.099 0.095 1.544 1.013 0.164 0.540
 Table 2 summarizes the performance and stability of all forecasting methods evaluated in this study, across both the M5 and VN1 datasets.
-For each method, the table reports four key metrics per dataset: RMSSE (Root Mean Squared Scaled Error) to evaluate point forecast accuracy, MQL (Multi-Quantile Loss) to assess probabilistic forecast performance, sMAPC (symmetric Mean Absolute Percentage Change) to compare the stability in point forecast settings, and MQC (Multi-Quantile Change) to understand the stability in probabilistic terms.
-Overall 111 The overall results are based on the baseline retraining scenario, $r=7$ for M5 and $r=1$ for VN1, which is commonly considered the standard in both theoretical literature and practical forecasting applications.
-, the models evaluated achieved better stability results on the M5 dataset than on VN1.
-Several factors may account for this difference, such as the larger dataset size and higher frequency of the M5 time series, the availability of rich external regressors (e.g., promotions, special events), and the existence of well-established benchmark hyperparameter settings, many of which were unavailable for VN1 at the time of model training.
-Moreover, in terms of point forecasting, machine learning models are much more stable than deep learning methods on both datasets, implying that the former are less sensitive to frequent updates.
+For each method, the table reports four key metrics per dataset: RMSSE (Root Mean Squared Scaled Error) to evaluate point forecast accuracy, SMQL (Scaled Multi-Quantile Loss) to assess probabilistic forecast performance, sMAPC (symmetric Mean Absolute Percentage Change) to compare the stability in point forecast settings, and SMQC (Scaled Multi-Quantile Change) to understand the stability in probabilistic terms.
+Overall, the models evaluated achieved better stability results on the M5 dataset than on VN1.
+Several factors may account for this difference, such as the larger dataset size and a higher frequency of the M5 time series, the availability of rich external regressors (e.g., promotions, special events), and the existence of well-established benchmark hyperparameter settings, many of which were unavailable for VN1 at the time of the analysis.
+Moreover, in terms of point forecasting, machine learning models are more stable than deep learning methods on both datasets, implying that the former are less sensitive to frequent updates.
 On the contrary, in terms of probabilistic forecasting, deep learning architectures show less instability, meaning that these approaches can be a better option when retraining is performed frequently.
-In general, ensembling seems to be a good solution to produce stable predictions.
-However, combining the most accurate models is often not sufficient to reduce the instability.
-Indeed, ensembles show increasing returns in stability as more models are added, as opposed to the well-known phenomenon in the ensemble literature, where accuracy improvements show diminishing returns from adding more models (Zanotti, 2025a).
-Interestingly, larger ensembles are usually more effective for stability purposes, implying that model diversity may be a more relevant criterion to create a forecast combinations, in contrast to model performance.
-This pattern is particularly true for probabilistic forecasting, but it is also evident in point predictions.
-For instance, in the VN1 dataset, the Ens4A and Ens5A combinations achieve good stability results simply because they contain both deep learning and machine learning models, extending the combination’s diversity.
-Indeed, even if the deep learning models’ instability is on average 60%, it is enough to add one single machine learning model to obtain an improvement of almost 45%.
+In general, ensembling seems to be a good solution to produce stable predictions when compared to base models.
+However, larger ensembles do not necessarily imply more stability, and it is not clear whether model diversity may have an impact on ensembles stability of predictions.
 Figure 1 shows the point forecast stability of each model along the different retraining scenarios for the M5 and the VN1 datasets.
 To allow comparisons, we display the results in relative terms with respect to the baseline scenarios, that is $r=7$ for M5 and $r=1$ for VN1.
-As expected, the sMAPC profiles are non-increasing functions of the retraining frequencies.
+As expected, the sMAPC profiles are non-increasing functions of the retraining scenario.
 The stability remains practically the same across update frequencies or even improves as the retraining period increases.
-Indeed, moving from high to low retraining frequencies, the stability of most global models improves compared to the baseline.
+Indeed, moving from a high to a low retraining scenario, the stability of most global models improves compared to the baseline 444 Some models, like LR and MLP, show slightly decreasing stability (higher sMAPC) with an increasing retraining period.
+This can be due to many factors (e.g.
+sensitivity to small data changes, lack of robustness to stale parameters), but further investigation is necessary.
+.
 In particular, for some models (i.e.
-CatBoost, LSTM, and TCN) the instability reduction is strong and consistent, reaching very low levels.
+CatBoost, LSTM, and TCN), the instability reduction is strong and consistent, reaching very low levels.
 These results imply that less frequent retraining may have a huge impact on the point forecast stability of global models.
@@
 In a similar fashion, Figure 4 summarizes the relative stability in a probabilistic forecasting setting.
-In this context, we observe that, for the M5 dataset, the stability (as measured by Multi Quantile Change) paths are constant across almost every retraining scenario, meaning that less frequent updates neither harm nor improve the probabilistic forecasting stability of most global models.
-For the VN1 dataset, instead, we observe an almost convex relationship between the stability and the retraining frequency.
-On average, models’ stability improves for low retraining scenarios and then starts deteriorating around $r=10$, implying that it is possible to avoid retraining up to 10 weeks without negatively impacting on the forecast stability.
-For both datasets, these results are confirmed by the statistical tests on MQC over the different retrain periods (see supplementary materials).
-Figure 4: MQC results for each method and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
+In this context, we observe that, for the M5 dataset, the stability (as measured by Scaled Multi Quantile Change) paths are constant across almost every retraining scenario, meaning that less frequent updates neither harm nor improve the probabilistic forecasting stability of most global models 555 There are a few strange spikes in the SMQC profiles of some models.
+These can be due to different factors (e.g.
+inappropriate hyperparameter values), but further investigation is necessary.
+.
+For the VN1 dataset, instead, we observe some differences among the models.
+Some of them show constant improvements in stability as the retraining scenario increases (i.e.
+NBEATS, NHITS, and CatBoost), while most of them do not exhibit any significant enhancements or deteriorations in the retraining intervals.
+For both datasets, these results are confirmed by the statistical tests on SMQC over the different retraining periods (see Supplementary materials).
+Figure 4: SMQC results for each method and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
 Overall, the results from Figures 1 and 4, supported by the Friedman-Nemenyi tests, indicate that reducing the retraining frequency of global models does not negatively impact (and may even enhance) forecast stability, for both point and probabilistic forecasts.
 When considered alongside the findings of Zanotti (2025b), this provides strong evidence against the practice of continuous retraining in global forecasting models.
-Lower retraining frequencies preserve both accuracy and stability, while offering substantial savings in computational resources.
-Figure 5 shows the evolution of point forecast stability, across different retraining scenarios for various ensemble configurations.
-The results on the M5 dataset indicate that reducing retraining frequency may negatively affect forecast stability.
+Lower retraining scenarios preserve both accuracy and stability, while offering substantial savings in computational resources.
+Figure 5 shows the evolution of point forecast stability, across different retraining scenarios for the various ensemble configurations.
+The results on the M5 dataset indicate that reducing retraining scenario may negatively affect forecast stability.
 However, this impact is below 1% for most retraining scenarios.
 Indeed, statistical tests (in Supplementary material) on the significance of these differences show that most retraining scenarios are not statistically different from each other, with periodic retraining strategies being at least as good as continuous retraining.
-The results on the VN1, instead, indicate that there exists a clear benefit in lowering the retraining frequency in terms of forecast stability.
-Notably, on the M5 dataset the smallest ensemble (Ens2A) consistently achieves the lowest sMAPC values, since it is obtained by combining XGBoost and LGBM that where both the most accurate and the most stable global forecasting models on that set of data.
-Nonetheless, larger ensembles (such as Ens4A and Ens5A) stability profiles tend to outperform smaller ones, especially at lower retraining frequencies, suggesting that model diversity plays a critical role in stabilizing ensemble forecasts when.
+The results on the VN1, instead, indicate that there exists a clear benefit in lowering the retraining scenario in terms of forecast stability.
+Notably, on the M5 dataset the smallest ensemble (Ens2A) consistently achieves the lowest sMAPC values, since it is obtained by combining XGBoost and LGBM, that where both the most accurate and the most stable global forecasting models on that set of data.
+On the contrary, larger ensembles (such as Ens4A and Ens5A) stability profiles tend to outperform smaller ones in the VN1 dataset, suggesting that when base models are less accurate, bigger ensembles may be necessary to improve forecast stability.
 Figure 6 further supports these observations in the probabilistic domain.
-First of all, probabilistic forecasts of ensembles are generally more stable than that of the base models across retraining frequencies, especially in M5, where MQC variations are within 1 percentage point.
-Moreover, in the VN1 dataset, the almost convex relationship between stability and the retraining frequency is flattened, because the variations are more restrained as the retraining period increases, and the decline in stability happens only from $r=26$ onwards.
-Then, larger ensembles provide consistently more stable probabilistic performance across most retraining scenarios, confirming the role of the diversity composition of ensembles even in the probabilistic setting.
-Most importantly, the effect of ensembling on the profiles of stability, as measured by sMAPC or MQC, is that it greatly reduces the variations across the different retraining frequencies, reducing both positive and negative stability changes.
+First of all, probabilistic forecasts of ensembles are generally more stable than those of the base models across retraining frequencies, especially in M5, where SMQC variations are within 1 percentage point.
+In the VN1 dataset, the stability improvement is evident as the models are retrained less frequently.
+Then, from a probabilistic perspective, all the different ensembles perform similarly in terms of stability of the predictions, without showing any relevant benefit of having larger ensembles with respect to smaller ones.
+Most importantly, the effect of ensembling on the profiles of stability, as measured by sMAPC or SMQC, is that it greatly reduces the variations across the different retraining scenarios, reducing both positive and negative stability changes.
 Figure 5: sMAPC results for each ensemble and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
-Figure 6: MQC results for each ensemble and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
-Figures 7 and 8 illustrate the fundamental trade-off between forecast accuracy and stability, comparing RMSSE with sMAPC and MQL with MQC, respectively.
-These plots reveal that high accuracy does not guarantee high stability.
-In particular, several deep learning models, achieve low RMSSE or MQL scores but exhibit high instability, especially in VN1, highlighting their sensitivity to changes in the training data and their volatility across forecast origins.
-Conversely, models like Random Forest, XGBoost, and LGBM strike a more favorable balance, maintaining relatively low RMSSE and sMAPC values.
-Nevertheless, it is the ensemble models, and especially those with larger size like Ens4A and Ens5A, that most consistently approach the optimality of stability and accuracy.
+Figure 6: SMQC results for each ensemble and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
+Figures 7 and 8 illustrate the fundamental trade-off between forecast accuracy and stability, comparing RMSSE with sMAPC and SMQL with SMQC, respectively.
+These plots reveal that high accuracy does not always guarantee high stability.
+In particular, several models achieve low RMSSE or SMQL scores but exhibit high instability, especially in VN1, highlighting their sensitivity to changes in the training data and their volatility across forecast origins.
+Nevertheless, it is the ensemble models that most consistently approach the optimality of stability and accuracy.
 In both the point and probabilistic domains, even very simple ensembles manage to control variance and mitigate instability without sacrificing much, if any, predictive performance.
-The overall picture that emerges from these figures reinforces the idea ensembling offers a practical and effective strategy to counteract forecast instability, particularly when constructed from a diverse pool of models.
-Figure 7: The accuracy-stability trade-off. RMSSE vs sMAPC for the M5 and VN1 datasets.
-Figure 8: The accuracy-stability trade-off. MQL vs MQC for the M5 and VN1 datasets.
-This set of results has clear practical implications: in operational forecasting systems, especially in domains such as retail and supply chain management, the use of ensembles combined with lower retraining frequencies may offer a robust forecasting process, balancing performance with reliability.
+The overall picture that emerges from these figures reinforces the idea ensembling offers a practical and effective strategy to counteract the forecast instability.
+Figure 7: The accuracy-stability trade-off. RMSSE vs sMAPC for the baseline scenario, $r=7$ for the M5 dataset, and $r=1$ for the VN1 dataset.
+Figure 8: The accuracy-stability trade-off. SMQL vs SMQC for the baseline scenario, $r=7$ for the M5 dataset, and $r=1$ for the VN1 dataset.
+This set of results has clear practical implications: in operational forecasting systems, especially in domains such as retail and supply chain management, the use of ensembles combined with lower retraining scenarios may offer a robust forecasting process, balancing performance with stability.
 5 Conclusions
-This study provides a comprehensive empirical assessment of forecast stability in global models, focusing on the effects of retraining frequency and ensembling across both point and probabilistic forecasting contexts.
-Our findings demonstrate that global forecasting models are able to produce very stable predictions and that they can achieve even more stability under reduced retraining frequencies, challenging the prevailing wisdom that frequent model updates are necessary to maintain forecast reliability.
-In terms of point forecasting, the results consistently show that retraining less frequently does not harm, and often improves, forecast stability.
-This trend is particularly evident in the M5 dataset, where many models exhibit monotonic or flat stability profiles as the retraining interval increases.
-Similar patterns, although more nuanced, are observed in the VN1 dataset.
-Probabilistic forecast stability, as measured by the newly defined Multi-Quantile Change (MQC), shows a slightly different behavior.
-While in the M5 dataset forecast stability remains largely unaffected by retraining frequency, in VN1 we observe a convex relationship, indicating that retraining up to every 10 weeks can preserve stability, but beyond that point the quality of probabilistic forecasts begins to deteriorate.
+This study provides a comprehensive empirical assessment of the forecast stability of global models, focusing on the effects of retraining across both point and probabilistic forecasting, in the context of retail demand.
+Our findings demonstrate that global forecasting models are able to produce very stable predictions, and that they can achieve even more stability under reduced retraining strategies, challenging the prevailing wisdom that frequent model updates are necessary to maintain forecast stability.
+In terms of point forecasting, the results consistently show that retraining less frequently does not harm and often improves forecast stability.
+This trend is evident in both the M5 and the VN1 datasets, where many models exhibit monotonic or flat stability profiles as the retraining interval increases.
+Probabilistic forecast stability, as measured by the newly defined Scaled Multi-Quantile Change (SMQC), shows a very similar behavior, indicating that the stability of probabilistic forecasts does not deteriorate as the models are retrained less frequently.
 These results are statistically validated via Friedman-Nemenyi tests and indicate that continuous retraining is often not the optimal strategy.
-Ensembling emerges as a powerful strategy to enhance forecast stability.
-While combining only the most accurate models offers moderate gains, the most substantial improvements in stability, both point and probabilistic, are achieved by ensembles that incorporate a diverse mix of forecasting algorithms.
-This effect is especially pronounced in the VN1 dataset, where larger and more heterogeneous ensembles significantly reduce forecast volatility across retraining scenarios.
-Notably, ensemble models not only improve stability metrics but also smooth the variation in stability across retraining frequencies, further reinforcing their role as a stabilizing force in global forecasting systems.
-These results suggest that model diversity should be prioritized over raw model performance when building ensembles.
+Ensembling can be considered a good strategy to enhance forecast stability.
+Notably, ensemble models not only improve stability metrics but also smooth the variation in stability across retraining scenarios, further reinforcing their role as a stabilizing force in global forecasting systems.
+However, from our analysis, it is not clear whether the size and the diversity of the ensemble may have an impact on stability, and more research is needed in this direction.
 The joint analysis of retraining and ensembling reveals that the two approaches act as complementary levers for stabilizing forecasts.
@@
 From a practical standpoint, these findings have several implications.
-First, businesses that rely on large-scale forecasting systems, such as those in retail, inventory management, and supply chain planning, can adopt lower retraining frequencies not only to conserve computational resources but also to increase operational stability, without compromising on forecast quality.
-Second, incorporating diverse ensembles into global forecasting frameworks provides an effective hedge against model instability, thereby increasing the reliability of forecasts used in decision-making processes.
+First, businesses that rely on large-scale forecasting systems, such as those in retail, inventory management, and supply chain planning can adopt lower retraining strategies not only to conserve computational resources, but also to increase operational stability, without compromising on forecast accuracy.
+Second, incorporating ensembles into global forecasting frameworks provides an effective hedge against model instability, thereby increasing the reliability of forecasts used in decision-making processes.
 Finally, these results support a shift in emphasis from pure accuracy to stability-aware model selection and evaluation, in those contexts that are more sensitive to forecast reviews.
@@
 While the analysis covers a broad range of global models and two state-of-the-art retail forecasting datasets, the generalizability of the findings to other domains or data types (e.g., macroeconomic indicators, financial series, or healthcare time series) remains an open question.
+Moreover, this work deliberately focuses on the stability of global models, leaving the analysis of stability in local models as an open avenue.
+A direct comparison between global and local approaches could therefore represent a valuable direction for future research.
 Additionally, the study focuses solely on full retraining schemes and does not explore incremental or online learning strategies, which may offer a different trade-off between accuracy and stability.
 Future research could extend this work by incorporating such adaptive methods, assessing alternative ensembling techniques (e.g., weighted or stacking ensembles), and introducing business-oriented stability metrics that link forecast instability directly to operational cost or risk.
-Further theoretical development of stability measures, particularly for probabilistic forecasts, may also contribute to a deeper understanding of how models behave across time and under varying data regimes.
-In conclusion, this work provides strong empirical evidence that frequent retraining is not necessary for maintaining stability in global forecasting models, and that ensembling, particularly when built on diverse model pools, is a highly effective strategy for achieving both stable and accurate forecasts.
+Furthermore, this study employed a single framework for generating probabilistic forecasts, namely Conformal Inference.
+While conformalized residuals provide a valid means for quantifying uncertainty, the resulting quantiles may primarily capture the stability of the conformal adjustment rather than that of the underlying predictive distributions.
+Future research could address this limitation by studying probabilistic stability in native probabilistic models, thereby enhancing the robustness and generalizability of the findings.
+Finally, further theoretical development of stability measures, particularly for probabilistic forecasts, may also contribute to a deeper understanding of how models behave across time and under varying data regimes.
+In conclusion, this work provides strong empirical evidence that frequent retraining is not necessary for maintaining stability in global forecasting models, and that ensembling is an effective strategy for achieving both stable and accurate forecasts.
 These findings contribute to a growing body of research promoting sustainable and robust forecasting practices in modern data-intensive environments.
 Supplementary material
-In this section, we provide tables and figures related to the empirical results of the M5 and VN1 datasets. Table 4 shows the composition of ensembles across the different datasets.
+In this section, we provide tables and figures related to the empirical results of the M5 and VN1 datasets.
+Table 3 shows the composition of ensembles across the different datasets.
+The Tables 4, 5, 7, and 6 show the forecast instability of the different models and ensembles along the examined retrain scenarios for the M5 daily dataset.
+The Tables 8, 9, 11, and 10 show the forecast instability of the different models and ensembles along the examined retrain scenarios for the VN1 weekly dataset.
+Figures 9, and 10 show the results of the Friedman-Nemenyi test in the context of probabilistic instability for the M5 and VN1 datasets.
+Figures 11, 12, 13, and 14 show the results of the Friedman-Nemenyi test in the context of both point and probabilistic instability for the different ensemble models across retrain scenarios and for each dataset.
 Table 3: Composition of ensembles for the M5 and VN1 datasets.
@@
 ENS5A MLP, NBEATSx, NHITS, RF, XGBoost
-The Tables 5, 6, 8, and 7 show the forecast instability of the different models and ensembles along the examined retrain scenarios for the M5 daily dataset.
 Method 7 14 21 30 60 90 120 150 180 364
@@
 Method 7 14 21 30 60 90 120 150 180 364
-LR 0.118 0.118 0.118 0.119 0.119 0.119 0.119 0.119 0.119 0.119
-XGBoost 0.115 0.115 0.115 0.116 0.115 0.115 0.115 0.116 0.115 0.114
-LGBM 0.123 0.122 0.123 0.124 0.123 0.123 0.123 0.124 0.123 0.122
-CatBoost 0.127 0.144 0.123 0.144 0.144 0.144 0.144 0.142 0.141 0.139
-MLP 0.109 0.109 0.110 0.110 0.111 0.111 0.112 0.111 0.110 0.110
-TCN 0.114 0.114 0.114 0.113 0.113 0.112 0.113 0.112 0.110 0.108
-NBEATSx 0.110 0.109 0.110 0.115 0.110 0.112 0.111 0.111 0.129 0.169
-NHITS 0.112 0.111 0.122 0.112 0.114 0.114 0.115 0.135 0.114 0.115
-Table 7: M5 MQC values for each method and retrain scenario combination.
+LR 0.110 0.110 0.110 0.111 0.111 0.111 0.112 0.112 0.112 0.113
+XGBoost 0.100 0.099 0.099 0.100 0.100 0.100 0.099 0.100 0.099 0.098
+LGBM 0.118 0.117 0.117 0.119 0.118 0.118 0.118 0.119 0.118 0.118
+CatBoost 0.156 0.157 0.154 0.157 0.157 0.157 0.157 0.156 0.156 0.155
+MLP 0.083 0.083 0.083 0.083 0.084 0.084 0.085 0.085 0.083 0.085
+TCN 0.083 0.083 0.083 0.082 0.082 0.082 0.082 0.082 0.080 0.080
+NBEATSx 0.084 0.084 0.084 0.088 0.085 0.085 0.085 0.086 0.088 0.105
+NHITS 0.087 0.085 0.101 0.086 0.087 0.088 0.088 0.091 0.088 0.089
+Table 6: M5 SMQC values for each method and retrain scenario combination.
 Method 7 14 21 30 60 90 120 150 180 364
-Ens2A 0.117 0.117 0.117 0.118 0.118 0.118 0.118 0.119 0.118 0.117
-Ens3A 0.117 0.117 0.116 0.118 0.118 0.118 0.117 0.118 0.117 0.117
-Ens4A 0.113 0.113 0.113 0.114 0.114 0.113 0.113 0.114 0.113 0.113
-Ens5A 0.111 0.111 0.111 0.112 0.111 0.111 0.111 0.111 0.110 0.110
-Table 8: M5 MQC values for each ensemble and retrain scenario combination.
-The Tables 9, 10, 12, and 11 show the forecast instability of the different models and ensembles along the examined retrain scenarios for the VN1 weekly dataset.
+Ens2A 0.108 0.108 0.108 0.108 0.108 0.108 0.108 0.108 0.108 0.107
+Ens3A 0.107 0.107 0.107 0.108 0.107 0.107 0.107 0.108 0.107 0.107
+Ens4A 0.099 0.099 0.099 0.101 0.100 0.100 0.099 0.100 0.099 0.099
+Ens5A 0.095 0.095 0.095 0.096 0.096 0.095 0.095 0.096 0.095 0.095
+Table 7: M5 SMQC values for each ensemble and retrain scenario combination.
 Method 1 2 3 4 6 8 10 13 26 52
@@
 Method 1 2 3 4 6 8 10 13 26 52
-LR 1.064 1.067 1.063 1.061 1.070 1.079 1.103 1.097 1.127 1.163
-RF 1.322 1.329 1.328 1.341 1.353 1.366 1.403 1.392 1.439 1.504
-XGBoost 1.101 1.072 1.080 1.079 1.083 1.097 1.084 1.135 1.175 1.254
-LGBM 1.282 1.179 1.235 1.161 1.165 1.173 1.199 1.277 1.266 1.379
-CatBoost 1.435 1.281 1.244 1.159 1.150 1.166 1.151 1.225 1.207 1.246
-MLP 0.844 0.876 0.887 0.883 0.889 0.891 0.876 0.933 0.903 0.969
-LSTM 1.020 1.015 1.008 1.004 1.005 0.993 0.986 0.987 0.972 0.930
-TCN 1.020 1.015 1.008 1.004 1.005 0.993 0.986 0.987 0.972 0.930
-NBEATSx 1.532 1.519 1.467 1.253 1.167 1.210 1.196 1.069 1.084 1.147
-NHITS 1.592 1.505 1.412 1.254 1.212 0.998 1.174 1.028 1.082 1.145
-Table 11: VN1 MQC values for each method and retrain scenario combination.
+LR 1.560 1.552 1.546 1.543 1.562 1.525 1.547 1.643 1.682 1.900
+RF 0.729 0.732 0.735 0.737 0.735 0.738 0.743 0.734 0.731 0.734
+XGBoost 0.629 0.598 0.600 0.584 0.585 0.575 0.556 0.613 0.599 0.589
+LGBM 1.484 1.433 1.457 1.471 1.502 1.349 1.459 1.394 1.593 1.445
+CatBoost 2.276 1.861 1.769 1.582 1.547 1.579 1.333 1.581 1.302 1.243
+MLP 0.219 0.281 0.278 0.278 0.282 0.275 0.266 0.285 0.267 0.283
+LSTM 0.242 0.238 0.237 0.235 0.233 0.229 0.227 0.224 0.214 0.187
+TCN 0.241 0.238 0.237 0.235 0.233 0.229 0.227 0.224 0.214 0.187
+NBEATSx 0.765 0.763 0.708 0.567 0.477 0.442 0.520 0.348 0.362 0.369
+NHITS 0.755 0.779 0.660 0.555 0.506 0.333 0.507 0.348 0.343 0.365
+Table 10: VN1 SMQC values for each method and retrain scenario combination.
 Method 7 14 21 30 60 90 120 150 180 364
-Ens2A 1.079 1.114 1.062 0.974 0.949 0.982 0.962 0.921 0.944 1.022
-Ens3A 1.200 1.192 1.148 1.024 0.990 0.973 0.995 1.005 0.970 1.034
-Ens4A 1.180 1.166 1.122 1.036 1.021 1.001 1.028 0.988 1.017 1.080
-Ens5A 1.153 1.137 1.107 1.033 1.022 1.000 1.031 1.006 1.029 1.091
-Table 12: VN1 MQC values for each ensemble and retrain scenario combination.
-Figures 9, and 10 show the results of the Friedman-Nemenyi test in the context of probabilistic instability for the M5 and VN1 datasets.
-Figure 9: M5 Friedman-Nemenyi test results based on MQC.
-Figure 10: VN1 Friedman-Nemenyi test results based on MQC.
-Figures 11, 12, 13, and 14 show the results of the Friedman-Nemenyi test in the context of both point and probabilistic instability for the different ensemble models across retrain scenarios and for each dataset.
+Ens2A 0.441 0.476 0.452 0.391 0.354 0.332 0.364 0.300 0.299 0.312
+Ens3A 0.539 0.562 0.507 0.430 0.389 0.324 0.398 0.312 0.309 0.322
+Ens4A 0.545 0.557 0.522 0.465 0.433 0.387 0.443 0.376 0.376 0.389
+Ens5A 0.540 0.545 0.519 0.472 0.448 0.410 0.453 0.407 0.405 0.416
+Table 11: VN1 SMQC values for each ensemble and retrain scenario combination.
+Figure 9: M5 Friedman-Nemenyi test results based on SMQC.
+Figure 10: VN1 Friedman-Nemenyi test results based on SMQC.
 Figure 11: M5 Friedman-Nemenyi test results on ensembles based on sMAPC.
-Figure 12: M5 Friedman-Nemenyi test results on ensembles based on MQC.
+Figure 12: M5 Friedman-Nemenyi test results on ensembles based on SMQC.
 Figure 13: VN1 Friedman-Nemenyi test results on ensembles based on sMAPC.
-Figure 14: VN1 Friedman-Nemenyi test results on ensembles based on MQC.
+Figure 14: VN1 Friedman-Nemenyi test results on ensembles based on SMQC.

### v3

@@
 Despite its importance, stability is often overlooked in favor of accuracy.
-In this study, we evaluate the stability of point and probabilistic forecasts across several retraining scenarios using two large retail demand datasets (M5 and VN1) and ten different global forecasting models.
-To analyze stability in the probabilistic setting, we propose a new model-agnostic, distribution-free, and scale-free metric that measures probabilistic instability: the Scaled Multi-Quantile Change (SMQC).
-Furthermore, we also evaluate the effects of retraining on various ensemble configurations based on forecast pooling.
-The results show that, compared to continuous retraining, less frequent retraining not only preserves but often improves forecast stability, challenging the need for continuous retraining.
-The study promotes a shift toward stability-aware forecasting practices, proposing a new tool to effectively evaluate forecast stability in probabilistic settings, and offering practical guidelines for building more robust prediction systems.
+In this study, we evaluate the stability of point and probabilistic forecasts across several retraining scenarios using three large forecasting datasets and ten different global forecasting models.
+To analyze stability in the probabilistic setting, we propose a new model-agnostic, distribution-free, and scale-free metric that measures probabilistic stability: the Scaled Multi-Quantile Change (SMQC).
+The results show that less frequent retraining not only preserves but often improves forecast stability, challenging the need for frequent retraining.
+Moreover, the study shows that accuracy and stability are not necessarily conflicting objectives when adopting a global modeling approach.
+The study promotes a shift toward stability-aware forecasting practices, proposing a new metric to evaluate forecast stability effectively in probabilistic settings, and offering practical guidelines for building more stable and sustainable forecasting systems.
 keywords:
-Time series, Demand forecasting, Forecasting competitions, Cross-learning, Global models, Forecast stability , Vertical stability , Machine learning, Deep learning, Conformal predictions
+Time series, Forecasting competitions, Cross-learning, Global models, Forecast stability , Vertical stability , Machine learning, Deep learning, Conformal predictions
+journal: International Journal of Forecasting
 1 Introduction
-In recent years, global forecasting models, those trained across multiple time series simultaneously, have emerged as a powerful alternative to traditional local approaches, particularly in large-scale applications such as retail demand forecasting.
+In recent years, global forecasting models, those trained across multiple time series simultaneously, have emerged as a powerful alternative to traditional local approaches, particularly in large-scale applications such as retail demand and energy forecasting.
 While their accuracy and efficiency have been widely studied, less attention has been paid to the stability of their forecasts over time.
@@
 As new information becomes available, updated forecasts may naturally differ from earlier versions, ideally improving in accuracy at the expense of stability.
-Conversely, overly stable forecasts might sacrifice accuracy by ignoring valuable new information.
+Conversely, overly stable forecasts might sacrifice accuracy by ignoring new, valuable information.
 Ensembling techniques, which integrate multiple forecasts from different models, have been proposed as a mechanism to concurrently reduce bias and variance (Wang et al., 2023).
-By aggregating diverse predictions, ensembling methods can alleviate the impact of individual model volatility and potentially achieve a more favorable equilibrium between accuracy and stability.
+By aggregating diverse predictions, ensembling methods can alleviate the impact of individual model volatility and potentially achieve a more favorable equilibrium between accuracy and stability, but at the expense of a severe increase in computational costs.
 Even if the trade-off between forecasting accuracy and stability presents a significant challenge for organizations seeking to balance predictive performance with operational stability, recent research suggests that high levels of stability and accuracy can be achieved at the same time ((Godahewa et al., 2025), (Van Belle et al., 2023)).
-Moreover, Zanotti (2025b) demonstrated that the retraining frequency does not harm the forecast accuracy of global models, meaning that models estimated using this forecasting approach retain the same level of performance even when updated less frequently.
-This has a strong potential implication on the forecast stability: if we can reduce the retraining frequency of global models, then we should obtain more stable predictions, because retraining leads to different model artifacts potentially producing stability disruptions.
+Moreover, Zanotti (2025) demonstrated that the retraining frequency does not harm the forecast accuracy of global models, suggesting that, when the data-generating process remains relatively stable, global models can preserve predictive performance even when updated less frequently.
+This finding raises a natural question regarding forecast stability.
+Since retraining inherently produces new model artifacts, the frequency with which models are retrained may influence the temporal consistency of forecasts.
+Whether less frequent retraining systematically improves, worsens, or leaves unchanged the stability of global forecasting models is therefore an important question that needs to be addressed.
 1.1 Research Question
-We aim to address the question ”Do global models for retail demand forecasting produce stable forecasts?”.
-Purposely, we study whether the global modeling approach produces stable forecasts, and we try to understand the effects of retraining on the forecast stability, that is whether avoiding re-estimation for every new observation damages the stability of global models.
-To address this question, we rely on the two most recent and comprehensive retail forecasting datasets: the M5 and VN1 competition data.
-To generally understand the stability (or instability) of the forecasting models, we consider ten distinct global forecasting methods (five from the traditional machine learning domain and five based on commonly used deep neural network architectures).
-Moreover, we examine stability across a range of retraining scenarios, from no retraining to continuous retraining, by exploring periodic strategies that broadly encompass the most practical and effective approaches.
-We also investigate the use of ensembling, or forecasting combinations, as a mean to obtain more stable forecasts.
-This approach can indeed be significant to mitigate forecast instability, reducing the model variance and bias.
+We aim to address the question ”How does the retraining frequency of global forecasting models affect the stability of their forecasts?”.
+Purposely, we study whether the global modeling approach produces stable forecasts, and we try to understand the effects of retraining on the forecast stability, that is, whether avoiding re-estimation for every new observation damages the stability of global models.
+To address this question, we rely on three of the most recent and comprehensive forecasting datasets: the M4, the M5, and the VN1 competition data.
+To generally understand the stability of the forecasting models, we consider ten distinct global forecasting methods (five from the traditional machine learning domain and five based on commonly used deep neural network architectures).
+We also compared the results with two well-known univariate local forecasting models.
+Moreover, we examine stability across a range of retraining scenarios, continuous retraining to no retraining, by exploring periodic strategies that broadly encompass the most practical and effective approaches.
 1.2 Contributions
@@
   • 3.
-    We suggest new metrics to evaluate the instability of probabilistic forecasts.
+    We suggest new metrics to evaluate the stability of probabilistic forecasts.
   • 4.
@@
 The concepts related to model update and retrain scenario are explained in 3.3, together with the evaluation strategy adopted, while the metrics used to assess the stability of models are shown in 3.4.
-In Section 4 we discuss the empirical findings of our study, including forecast stability and ensembling on the different scenarios.
+In Section 4 we discuss the empirical findings of our study, including forecast accuracy and stability on the different scenarios.
 Finally, Section 5 contains our summary and conclusions.
 2 Related works
+2.1 Global modeling
 The cross-learning approach has seen substantial development in recent years.
 Today, most time series forecasting studies include at least a benchmark comparison involving global models (GMs), underscoring their growing importance in the field.
-Following Zanotti (2025b), we define $\mathcal{Y}$ as the set of all available time series in a dataset, where each $Y_{i}\in\mathcal{Y}$ represents an individual time series.
+We define $\mathcal{Y}$ as the set of all available time series in a dataset, where each $Y_{i}\in\mathcal{Y}$ represents an individual time series.
 Let $\mathcal{F}$ denote the set of possible predictive functions, with each $F\in\mathcal{F}$ corresponding to a specific forecasting model.
 Without loss of generality, we assume that all necessary information for prediction is contained within $\mathcal{Y}$.
-Under the local modeling paradigm, forecasts for a horizon $h$ are generated by training a separate model for each individual time series.
+Under the local modeling paradigm (LMs), forecasts for a horizon $h$ are generated by training a separate model for each individual time series.
 This implies that each series $Y_{i}$ is associated with its own model, characterized by its own set of parameter values.
@@
 (2022) explored the conditions under which global models are competitive.
-Additionally, Montero-Manso & Hyndman (2021) and (Montero-Manso, 2023) provided theoretical support showing that GMs can match or surpass the accuracy of local models, with lower complexity and without assuming data similarity.
+Additionally, Montero-Manso and Hyndman (2021) and (Montero-Manso, 2023) provided theoretical support showing that GMs can match or surpass the accuracy of LMs, with lower complexity and without assuming data similarity.
 Global models have proven to be the most accurate method in several forecasting domains, including retail demand (Spiliotis et al.
 (2022), Bandara et al.
-(2019), Juan R Trapero & Fildes (2015)), electricity demand (Buonanno et al., 2022), water demand (Groß & Hans, 2024), gas consumption (Gaweł & Paliński, 2024), and crop production (Ibañez & Monterola, 2023).
-Their effectiveness was particularly evident during the M5 competition (Makridakis et al., 2022a), where tree-based models employing cross-learning ranked among the top-performing solutions (Januschowski et al., 2022).
+(2019), Juan R Trapero and Fildes (2015)), electricity demand (Buonanno et al., 2022), water demand (Groß and Hans, 2024), gas consumption (Gaweł and Paliński, 2024), and crop production (Ibañez and Monterola, 2023).
+Their effectiveness was particularly evident during the M5 competition (Makridakis et al., 2022b), where tree-based models employing cross-learning ranked among the top-performing solutions (Januschowski et al., 2022).
 To further enhance GM performance, several strategies such as clustering (Godahewa et al.
@@
 Moreover, new machine learning (Godahewa et al., 2023) and deep learning (Oreshkin et al., 2020) architectures have been specifically designed to support cross-learning.
-Recently, research has begun focusing on improving GMs’ ability to capture local patterns (Sen et al., 2019) and enhance their interpretability (Rajapaksha et al., 2023).
+Recently, research has begun focusing on improving GMs’ ability to capture local patterns (Sen et al., 2019) and enhance their explainability (Rajapaksha et al., 2023).
+2.2 Forecast stability
 From a forecasting stability perspective, the literature is impoverished.
@@
 (2023), instead, extended an existing deep learning architecture (NBEATS) to optimize forecasts from both a traditional forecast accuracy perspective, as well as a forecast stability perspective, directly including an instability component into the loss function of the model.
-There is also active research on forecast stability within judgmental forecasting ground (Fildes & Goodwin, 2021).
-However, almost all the literature related to forecast stability is focused on point prediction stability only, possibly because it is not clear how to measure instability in terms of probabilistic forecasting, and maybe also because most machine learning and deep learning methods do not directly output probabilistic predictions (Makridakis et al., 2022c).
+There is also active research on forecast stability within judgmental forecasting ground (Fildes and Goodwin, 2021).
+However, almost all the literature related to forecast stability is focused on point prediction stability only.
+This is not due to the absence of probabilistic distance measures, nor to the lack of probabilistic forecasting techniques (Makridakis et al., 2022a).
+Rather, probabilistic stability has received limited attention because existing evaluation metrics are typically designed to compare forecasts to realizations or to assess calibration and sharpness, rather than to quantify temporal consistency across successive forecast origins.
 Nevertheless, in many forecasting applications, such as supply chain management, it is crucial to generate and evaluate predictions probabilistically, whether through prediction intervals, quantiles, or full predictive distributions (Fildes et al., 2022).
@@
 Therefore, the stability evaluation of probabilistic forecasts is an open question, especially in a model-agnotic way.
-Lukily, among the methods developed for uncertainty quantification, Vovk et al.
+Luckily, among the methods developed for uncertainty quantification, Vovk et al.
 (2005) introduced Conformal Inference, which is a model-agnostic framework that offers valid uncertainty estimates and can also be applied in time series forecasting settings (Stankeviciute et al., 2021).
-In the context of model retraining and updating strategies, the most comprehensive work in time series forecasting is by Zanotti (2025b), who showed the positive effects of reducing retraining of global models on both accuracy and forecasting costs.
-The seminal work of Spiliotis & Petropoulos (2024) extensively investigated the impact of various retraining scenarios and parameter update methods on model performance, focusing on the exponential smoothing family of models, following the traditional local modeling approach.
-(Huber & Stuckenschmidt, 2020) touched on retraining within the retail demand domain, but the study was restricted to a small set of models, limited retraining strategies, and a proprietary daily dataset.
+2.3 Model retraining
+In the context of model retraining and updating strategies, the most comprehensive work in time series forecasting is by Zanotti (2025), who showed the positive effects of reducing retraining of global models on both accuracy and forecasting costs.
+The seminal work of Spiliotis and Petropoulos (2024) extensively investigated the impact of various retraining scenarios and parameter update methods on model performance, focusing on the exponential smoothing family of models, following the traditional local modeling approach.
+(Huber and Stuckenschmidt, 2020) touched on retraining within the retail demand domain, but the study was restricted to a small set of models, limited retraining strategies, and a proprietary daily dataset.
 While these findings are encouraging, there has been little direct investigation into whether global models specifically can retain stability with less frequent updates.
@@
 This section presents the empirical analysis carried out to investigate the stability of global models and to determine whether less frequent retraining scenarios can yield stability outcomes comparable to those of the baseline scenario, which involves the most frequent retraining.
-We begin by describing the datasets used in our experiments, followed by an overview of the machine learning, deep learning, and ensemble models employed.
+We begin by describing the datasets used in our experiments, followed by an overview of the machine learning and deep learning models employed.
 Finally, we detail the instability measures, the various retraining scenarios considered, and the evaluation strategy applied to assess forecast performance.
 3.1 Datasets
-For our experiments, we employed two retail forecasting datasets: the M5 and the VN1 competition datasets.
-The M5 competition, part of the well-known M-competitions series organized by Spyros Makridakis and colleagues, aimed to benchmark forecasting methods in the context of retail demand (Makridakis et al., 2022b).
-The M5 dataset (Howard & Makridakis, 2020) includes 30,490 daily time series representing unit sales of Walmart products across three categories (Food, Hobbies, and Household) sold in ten stores located in California, Texas, and Wisconsin.
+For our experiments, we employed three forecasting datasets: the M4, the M5, and the VN1 competition datasets.
+The M4 competition is part of the well-known M-competitions series organized in 2018 by Spyros Makridakis and colleagues Petropoulos and Makridakis (2020).
+It includes real-world time series drawn from diverse domains such as finance, economics, demographics, industry, and other indicators, and spans multiple frequencies.
+In our study, we focused on the 4,227 daily time series characterized by substantial variability in length, seasonality, and noise structure.
+The heterogeneous nature of the series makes it particularly well-suited for studying general retraining behaviors that are not domain-specific.
+The M5 competition, also part of the M-competitions series, was organized to benchmark forecasting methods in the context of retail demand (Makridakis et al., 2022c).
+The M5 dataset (Howard and Makridakis, 2020) includes 30,490 daily time series representing unit sales of Walmart products across three categories (Food, Hobbies, and Household) sold in ten stores located in California, Texas, and Wisconsin.
 The data span from 2011 to 2016 and are characterized by high intermittency and a hierarchical structure, enabling forecasts at multiple aggregation levels (e.g., SKU, product category, store, and state).
@@
 As far as we are aware, our study is among the first to evaluate forecasting models on this dataset.
-Together, these two datasets offer the most recent and comprehensive time series collections related to retail demand, enabling good generalizability of our findings.
-In both cases, our analysis focused on the most disaggregated level of the data (i.e., SKUs), where the potential advantages of retraining less frequently are likely to be most pronounced.
-To ensure consistency with the evaluation setup described in Section 3.3, we applied a filtering criterion: for the daily data (M5), only time series with at least two years of observations (730 data points) were retained; for the weekly data (VN1), we considered only those series with at least three years of data (157 observations).
+Together, the M5 and the VN1 datasets offer the most recent and comprehensive time series collections related to retail demand, enabling good generalizability of our findings also to such domain.
 3.2 Forecasting models
-In this section, we present an overview of the global models used in our experiments.
-In our study, we focused exclusively on analyzing the stability of global methods, as our primary goal was to assess whether this modeling approach, unlike the traditional local one, can maintain performance with less frequent retraining 111 The analysis of the stability of local models was out of the scope of our study.
+In our study, we focused on analyzing the stability of global methods, as our primary goal is to assess whether this modeling approach, unlike the traditional local one, can maintain performance with less frequent retraining.
+Nevertheless, we also compared the results with two popular local models, that is the Exponential Smoothing (ETS) (Hyndman et al., 2002) and the Autoregressive Integrated Moving Average (ARIMA) (Hyndman and Khandakar, 2008) methods in their automatic versions  [^1] A comprehensive analysis of retraining effects on the stability of local models was out of the scope of this study.
 .
-Cross-learning has become the standard in many industries dealing with large-scale time series data, particularly in retail demand forecasting, where regular forecasts for thousands of SKUs are required (Januschowski et al., 2020).
-Therefore, to conduct a comprehensive evaluation of the stability of global models, we incorporated both traditional machine learning algorithms and state-of-the-art deep learning architectures.
+To conduct a comprehensive evaluation of the stability of global models, we incorporated both traditional machine learning algorithms and state-of-the-art deep learning architectures.
 The selected models are well-established in the time series forecasting literature and represent a range of methodological paradigms, allowing for a broad and informative comparison.
@@
 While machine learning models are easier to train compared to deep learning alternatives, they often rely heavily on high-quality feature engineering (Januschowski et al., 2022).
-For our experiments, we adopted simplified versions of the feature engineering pipelines used in the top-performing M5 and VN1 solutions.
+For our experiments, we adopted simplified versions of the feature engineering pipelines used in the top-performing M4, M5 and VN1 solutions.
 We constructed time series features such as lags, rolling means, expanding means, and calendar features (e.g., year, month, week, day of the week).
-We also included static metadata such as store, product, category, and location identifiers.
+For the M5 and VN1 datasets, we also included static metadata such as store, product, category, and location identifiers.
 For the M5 dataset, additional external features like special events were also incorporated.
-Model hyperparameters were selected based on configurations used in top-performing competition solutions when available; otherwise, we relied on the recommended defaults from the respective libraries.
+Model hyperparameters were selected based on configurations used in top-performing competition solutions when available; otherwise, we relied on the recommended defaults from the respective automatic libraries.
 We compared five machine learning models for time series forecasting.
-Linear Regression (LR) is a classical statistical model effective with appropriate feature engineering, often used as a benchmark in global model performance ((Montero-Manso & Hyndman, 2021), (Godahewa et al., 2021b)).
+Linear Regression (LR) is a classical statistical model effective with appropriate feature engineering, often used as a benchmark in global model performance ((Montero-Manso and Hyndman, 2021), (Godahewa et al., 2021b)).
 Random Forest (RF), an ensemble learning method, captures non-linear patterns and is robust to overfitting, making it suitable for demand forecasting ((Breiman, 2001), (Januschowski et al., 2022)).
-Extreme Gradient Boosting (XGBoost) and Light Gradient Boosting Machine (LGBM) are gradient-boosted models known for speed and accuracy, with LGBM excelling in large, high-dimensional datasets ((Chen & Guestrin, 2016), (Ke et al., 2017), Makridakis et al.
-(2022a)).
+Extreme Gradient Boosting (XGBoost) and Light Gradient Boosting Machine (LGBM) are gradient-boosted models known for speed and accuracy, with LGBM excelling in large, high-dimensional datasets ((Chen and Guestrin, 2016), (Ke et al., 2017), Makridakis et al.
+(2022b)).
 Categorical Boosting (CatBoost) specializes in handling categorical data with minimal preprocessing (Prokhorenkova et al., 2018).
@@
 In our implementation, we followed competition-proven practices by providing only static, calendar, and external covariates, together with historical values, as inputs.
-Top solutions’ guidelines were adopted for setting the model hyperparameters when available, and otherwise, the recommended defaults from the respective libraries were used.
+Top solutions’ guidelines were adopted for setting the model hyperparameters when available, and otherwise, the recommended defaults from the respective automatic libraries were used.
 We compared five deep learning models for time series forecasting.
-Multi-Layer Perceptron (MLP) is a versatile, efficient neural network used in many time series forecasting tasks (Rosenblatt, 1958), while Recurrent Neural Networks (RNNs), particularly LSTMs and GRUs, capture temporal dependencies in sequential data ((Cho et al., 2014), (Hochreiter & Schmidhuber, 1997)).
+Multi-Layer Perceptron (MLP) is a versatile, efficient neural network used in many time series forecasting tasks (Rosenblatt, 1958), while Recurrent Neural Networks (RNNs), particularly LSTMs and GRUs, capture temporal dependencies in sequential data ((Cho et al., 2014), (Hochreiter and Schmidhuber, 1997)).
 Temporal Convolutional Networks (TCN) offer an alternative by using causal convolutions to capture long-range dependencies (Van den Oord et al., 2016).
 Neural Basis Expansion Analysis for Time Series (NBEATS) and Neural Hierarchical Interpolation for Time Series (NHITS) are state-of-the-art models for time series forecasting, with NBEATS offering interpretable trend and seasonality components (Oreshkin et al., 2020), and NHITS improving upon NBEATS with hierarchical interpolation mechanisms (Challu et al., 2022).
-Ensemble learning combines the predictions of multiple base models to improve forecast accuracy, particularly when individual models capture complementary patterns in the data.
-In the context of global forecasting, ensembles are especially valuable for mitigating the risk of model instability and overfitting, which can occur when relying on a single method (Wang et al., 2023).
-Following Zanotti (2025a), we estimated four different ensemble models, adopting a simple averaging strategy, which is widely used for its ease of implementation and effectiveness in improving forecast robustness (Claeskens et al., 2016).
-Indeed, simple averaging can be used for both point and probabilistic forecasting, where equally weighting quantiles has been shown to yield robust results Busetti (2017).
-To evaluate the potential benefits of ensembling, we simulated a range of ensemble models to assess how forecast stability across the full test window would have differed if combinations had been employed in place of individual base models.
-Specifically, we constructed simple mean ensembles by aggregating the forecasts of subsets of global models, selecting only those base learners that achieved the highest individual accuracy.
-This design leverages the strengths of top-performing models while aiming to mitigate their individual weaknesses through averaging.
-Ensembles of increasing size were considered, combining the top 2, 3, 4, and 5 models according to each out-of-sample performance.
-Such selective pooling is essentially a variant of model pooling in the sense of Kourentzes et al.
-(2019), where only a subset of candidate models contributes to the final forecast.
-Nevertheless, to maintain a balance between potential improvements and the diminishing returns or additional complexity typically associated with larger ensembles (Wang et al., 2023), we limited the maximum ensemble size to five models (out of the eight or ten available, depending on the dataset).
-The ensemble models were finally validated using the same rolling origin evaluation framework described in Section 3.3.
-By incorporating ensemble learning into our analysis, we aimed to study whether combining global models could further enhance forecasting stability under different retraining scenarios.
-All global models were implemented in Python using Nixtla’s framework (Nixtla, 2022).
-Specifically, the mlforecast library was used to train the machine learning models, while neuralforecast was employed for efficiently training the deep learning models.
+All models were implemented in Python using Nixtla’s framework (Nixtla, 2022).
+Specifically, the statsforecast library was adopted to train univariate statistical models, the mlforecast library was used to train the global machine learning models, while neuralforecast was employed for efficiently training the global deep learning models.
+For this study, we adopted a cloud computing machine NC6s_v3 hosted on Microsoft Azure, with Linux Ubuntu 24 operating system, 1 Graphical Processing Unit, 6 Computing Processing Units, 112GB of memory.
 3.3 Evaluation strategy
 Out-of-sample evaluation is a cornerstone of time series forecasting, providing a way to test a model’s generalization ability beyond the training period, an essential step given that future data may diverge from historical patterns due to structural breaks, level shifts, or unexpected shocks (Tashman, 2000).
-Among the various evaluation strategies, the rolling origin evaluation has emerged as the most widely accepted and rigorous method (Bergmeir & Benítez, 2012).
+Among the various evaluation strategies, the rolling origin evaluation has emerged as the most widely accepted and rigorous method (Bergmeir and Benítez, 2012).
 It respects the chronological nature of time series data while enabling repeated assessments across multiple forecast origins.
@@
 Evaluation metrics (see Section 3.4) are averaged over all iterations to provide a robust estimate of forecasting accuracy.
-Compared to fixed origin evaluations, rolling origin offers a more nuanced view of a model’s robustness by exposing it to a range of temporal conditions (including seasonal patterns, trends, and potential anomalies) that may not be captured in a single evaluation window (Bergmeir & Benítez, 2012).
-This iterative approach reduces the risk of overfitting to one specific train-test split, and is especially beneficial in operational contexts like retail, logistics, or finance, where forecasts are regularly updated in response to incoming data.
+Compared to fixed origin evaluations, rolling origin offers a more nuanced view of a model’s robustness by exposing it to a range of temporal conditions (including seasonal patterns, trends, and potential anomalies) that may not be captured in a single evaluation window (Bergmeir and Benítez, 2012).
+This iterative approach reduces the risk of overfitting to one specific train-test split, and is especially beneficial in operational contexts like retail, energy, logistics, or finance, where forecasts are regularly updated in response to incoming data.
 It also aligns well with real-world decision-making processes that require forecasts to adapt over time.
@@
 The expanding window strategy is particularly suited for short time series, as it maximizes the amount of historical data available at each iteration.
-This approach is widely used in applied settings (Petropoulos & et al., 2022), and it is the one we adopted in our study, especially given the relatively short length of the weekly VN1 series.
+This approach is widely used in applied settings (Petropoulos and et al., 2022), and it is the one we adopted in our study, especially given the relatively short length of the weekly VN1 series.
 The Table 1 outlines the key parameters used in our experiments, including a test set spanning at least one full year, to mitigate seasonal bias, horizons aligned with typical business use cases, and a step size of one to maximize evaluation frequency.
@@
 Table 1: Retraining set, train and test size, frequency, forecast horizon, and step size for each dataset. The setup is based on the respective time series frequency.
-M5 VN1
+M4 / M5 VN1
 Frequency ($f$) daily (7) weekly (52)
@@
 Retrain set ($r$) {7, 14, 21, 30, 60, 90, 120, 150, 180, 364} {1, 2, 3, 4, 6, 8, 10, 13, 26, 52}
-Furthermore, to investigate the impact of retraining on forecasting stability, we examined various retraining scenarios, or retraining windows, as defined in Zanotti (2025b).
-Each scenario $r$ indicates the number of new observations after which the model is retrained.
-These retraining scenarios are tailored to the data frequency, daily for M5 and weekly for VN1, since frequency dictates both the forecast horizon and business review cycles.
+Furthermore, to investigate the impact of retraining on forecasting stability, we examined various retraining scenarios, or retraining windows.
+As defined in Zanotti (2025), each scenario $r$ indicates the number of new observations after which the model is retrained.
+These retraining scenarios are tailored to the data frequency, daily for M4 and M5, and weekly for VN1, since frequency dictates both the forecast horizon and business review cycles.
 For instance, in the daily case, $r=7$ reflects weekly updates.
-Across all scenarios, we trained global models on aligned datasets, updating the model either completely at each retraining step or not at all.
+Across all scenarios, we trained global models on each dataset, updating the model either completely at each retraining step or not at all.
 We excluded hyperparameter tuning due to its high cost and marginal expected benefit.
-Following Zanotti (2025b), we treated $r=1$ (or $r=7$ for daily data) as the accuracy benchmark and $r=T$ as the no-retraining baseline, with intermediate values representing periodic retraining strategies.
+We treated $r=1$ (or $r=7$ for daily data) as the accuracy benchmark and $r=T$ as the no-retraining baseline, with intermediate values representing periodic retraining strategies.
 3.4 Evaluation metrics
+3.4.1 Performance metrics
+Evaluating point forecast accuracy in time series is a debated topic, with no consensus on the best metric (Hewamalage et al., 2023).
+In our context, characterized by intermittent data, metrics based on absolute or percentage errors are suboptimal (Kolassa, 2020), and due to varying scales across series, a scaled accuracy metric should be preferred.
+Hence, we adopted the Root Mean Squared Scaled Error (RMSSE) (Hyndman and Koehler, 2006),
+$\text{RMSSE}=\sqrt{\frac{\frac{1}{h}\sum_{t=n+1}^{n+h}(y_{t}-\hat{y}_{t})^{2}}{\frac{1}{n-s}\sum_{t=s+1}^{n}(y_{t}-y_{t-s})^{2}}}\mbox{.}$ (3)
+which compares the model’s squared error to that of a seasonal naive forecast, and it was the official accuracy metric in the M5 competition (with $s=1$) (Makridakis et al., 2022c).
+Lower values indicate better performance.
+To assess probabilistic accuracy, we evaluated the quantiles produced using a scaled version of the Multi-Quantile Loss.
+The Scaled Quantile Loss, also known as (Scaled) Pinball Loss, and the Scaled Multi-Quantile Loss are defined as:
+$\text{SQL}(\alpha)=\frac{\frac{1}{h}\sum_{t=n+1}^{n+h}\Big(\alpha\cdot(y_{t}-\hat{q}_{t}^{(\alpha)})\cdot\mathbb{I}_{\hat{q}_{t}^{(\alpha)}\leq y_{t}}+(1-\alpha)\cdot(\hat{q}_{t}^{(\alpha)}-y_{t})\cdot\mathbb{I}_{\hat{q}_{t}^{(\alpha)}>y_{t}}\Big)}{\frac{1}{n-s}\sum_{t=s+1}^{n}|y_{t}-y_{t-s}|}\mbox{,}$ (4)
+$\text{SMQL}=\frac{1}{|\mathcal{Q}|}\sum_{\alpha\in\mathcal{Q}}\text{SQL}(\alpha)\mbox{,}$ (5)
+These proper scoring rules assess forecast distribution accuracy (Kolassa, 2016).
+Moreover, the SMQL was the official metric of the M5 Uncertainty competition (weighted, and with $s=1$) (Makridakis et al., 2022a).
+We considered a total of 14 different quantiles of probability levels $\alpha\in$ {0.005, 0.025, 0.050, 0.100, 0.150, 0.200, 0.025, 0.750, 0.800, 0.850, 0.900, 0.950, 0.975, 0.995}.
+Therefore, we analyzed the 50% 60%, 70%, 80%, 90%, 95%, and 99% central prediction intervals.
+The 50% and 60% central prediction intervals provide a good description of the center of the forecast distribution, while the 90%, 95%, and 99% give more information about its tails.
+These quantiles provide sufficient information about the uncertainty of the forecasts and allow for the effective description of the whole distribution.
+To ensure reliable quantile estimates, conformal prediction intervals were computed on a validation set at least twice the forecast horizon, which constrained the number of series used  [^2] We used a validation set four times larger than the forecast horizon for the M4 and M5 datasets, and twice the forecast horizon for the VN1 dataset.
+This limited the number of time series from 4,227 to 2,557 for the M4 dataset, and from 30,490 to 28,298 for the M5 dataset.
+.
+We also normalized each evaluation metric relative to the baseline retraining scenario, defined by the dataset frequency, to enable consistent comparison across models and retraining periods.
+To statistically validate our findings, we applied the Friedman-Nemenyi test (Demšar, 2006).
 3.4.2 Stability metrics
@@
 The instability across different pairs of consecutive forecasting origins can be obtained by a simple average of the sMAPC values.
-Lower values imply less unstable predictions.
+Lower values imply more stable predictions.
 Beyond point forecasts, our study also emphasizes the probabilistic stability of the models under different retraining strategies.
 Since most machine learning and deep learning models do not natively produce full predictive distributions, we employed Conformal Inference to construct prediction intervals around point forecasts.
-Its advantages (validity guarantees, minimal assumptions, computational simplicity, and effectiveness with limited data) make it especially suitable for our global forecasting setup, where different models are compared across heterogeneous datasets and varying retraining scenarios 222 One may prefer, instead, to train models using quantile losses to directly yield quantile forecasts, or to use inherently probabilistic models.
-These are also valid ways of modeling uncertainty, as demonstrated by the M5 Uncertainty competition (Makridakis et al., 2022c; Lainder & Wolfinger, 2022).
+Its advantages (validity guarantees, minimal assumptions, computational simplicity, and effectiveness with limited data) make it especially suitable for our global forecasting setup, where different models are compared across heterogeneous datasets and varying retraining scenarios  [^3] One may prefer, instead, to train models using quantile losses to directly yield quantile forecasts, or to use inherently probabilistic models.
+These are also valid ways of modeling uncertainty, as demonstrated by the M5 Uncertainty competition (Makridakis et al., 2022a; Lainder and Wolfinger, 2022).
 However, comparing different probabilistic modeling frameworks is out of the scope of our analysis, and it may be an area of future research.
@@
 Given two forecast origins, $n$ and n1n-1, QC quantifies the average adjustment in the predicted quantiles across the forecast horizon $h$.
-The formula mirrors the Quantile Loss (Pinball Loss) but replaces the actual observations $y_{t}$, with previous quantile forecasts $\hat{q}_{t,n-1}^{(\alpha)}$, thereby shifting the focus from forecast accuracy to forecast stability.
+The formula mirrors the Quantile Loss but replaces the actual observations $y_{t}$, with previous quantile forecasts $\hat{q}_{t,n-1}^{(\alpha)}$, thereby shifting the focus from forecast accuracy to forecast stability.
 The intuition is straightforward: if the forecasts are stable across time (i.e., $\hat{q}_{t,n}^{(\alpha)}\approx\hat{q}_{t,n-1}^{(\alpha)}$), then QC will be small.
@@
 To further contextualize the proposed metric, it is important to compare it with existing distributional distance metrics such as the Kullback–Leibler ($KL$) divergence and the Wasserstein distance ($W_{2}$).
-KL divergence measures the relative entropy between two probability distributions and is particularly sensitive to differences in the tails (Kullback & Leibler, 1951).
+KL divergence measures the relative entropy between two probability distributions and is particularly sensitive to differences in the tails (Kullback and Leibler, 1951).
 However, it requires access to full predictive densities, which can limit its use in forecasting applications where outputs are often expressed only as quantiles.
-Wasserstein distance ($W_{2}$), by contrast, quantifies the minimal cost of transforming one distribution into another and effectively captures both location and shape differences, but it again relies on full distributions (Peyré & Cuturi, 2019).
-The MQC metric addresses a complementary objective: it evaluates the stability of probabilistic forecasts directly at the quantile level, without the need of full predictive densities.
+Wasserstein distance ($W_{2}$), by contrast, quantifies the minimal cost of transforming one distribution into another and effectively captures both location and shape differences, but it usually relies on full distributions (although it can be approximated through discretization) (Peyré and Cuturi, 2019).
+The MQC metric addresses a complementary objective: it evaluates the stability of probabilistic forecasts directly at the quantile level, without the need for full predictive densities.
 Since it is based solely on a finite set of quantiles rather than the full predictive distribution, MQC does not capture all aspects of the distributional change, potentially missing more subtle shifts in shape, skewness, or variance that occur between those quantiles.
 However, in practical forecasting applications, especially in retail and supply chain domains where decisions are often made based on specific quantile levels (e.g., high quantiles for safety stock), this limitation is mitigated by the fact that the selected quantiles are usually those of greatest operational interest.
-Additionally, given that it is based on the pinball loss, the MQC is inherently asymmetric, reflecting the fact that quantile forecasts are directional.
-For example, an unstable 90% quantile forecast may have stronger implications on inventory or risk management than a similarly unstable 60% quantile forecast.
-This implies that MQC emphasizes directional changes when measuring stability, in contrast to symmetric distances (like $W_{2}$), which treat deviations identically regardless of the quantile’s position.
 Importantly, MQC is both model-agnostic and distribution-free, meaning that it can be applied consistently across different forecasting approaches without requiring strong assumptions about the underlying predictive distributions.
 Moreover, MQC is computationally lightweight compared to the Wasserstein distance or the KL divergence, which makes it suitable for large-scale applications involving thousands of time series.
-We considered a total of 14 different quantiles of probability levels $\alpha\in$ {0.005, 0.025, 0.050, 0.100, 0.150, 0.200, 0.025, 0.750, 0.800, 0.850, 0.900, 0.950, 0.975, 0.995}.
-Therefore, we analyzed the 50% 60%, 70%, 80%, 90%, 95%, and 99% central prediction intervals.
-The 50% and 60% central prediction intervals provide a good description of the center of the forecast distribution, while the 90%, 95%, and 99% give more information about its tails, essential to determine the safety stock levels in retail demand forecasting (Barrow & Kourentzes, 2016).
-These quantiles provide sufficient information about the uncertainty of the forecasts and allow for the effective description of the whole distribution.
-To ensure reliable quantile estimates, conformal prediction intervals were computed on a validation set at least twice the forecast horizon, which constrained the number of time series used 333 We used a validation set four times larger than the forecast horizon for the M5 dataset and twice the forecast horizon for the VN1 dataset.
-This limited the number of time series from 30,490 to 28,298 for the M5 dataset.
-.
-We also normalized each evaluation metric relative to the baseline retraining scenario, defined by the dataset frequency, to enable consistent comparison across models and retraining periods.
-To statistically validate our findings, we applied the Friedman-Nemenyi test for multiple comparisons (Demšar, 2006).
-3.4.2 Performance metrics
-Evaluating point forecast accuracy in time series is a debated topic, with no consensus on the best metric (Hewamalage et al., 2023).
-In our context of SKU-level demand forecasting, characterized by intermittent data, metrics based on absolute or percentage errors are suboptimal (Kolassa, 2020), and due to varying scales across series, a scaled accuracy metric has to be preferred.
-Hence, we adopted the Root Mean Squared Scaled Error (RMSSE) (Hyndman & Koehler, 2006),
-$\text{RMSSE}=\sqrt{\frac{\frac{1}{h}\sum_{t=n+1}^{n+h}(y_{t}-\hat{y}_{t})^{2}}{\frac{1}{n-s}\sum_{t=s+1}^{n}(y_{t}-y_{t-s})^{2}}}\mbox{.}$ (8)
-which compares the model’s squared error to that of a seasonal naive forecast, and it was the official accuracy metric in the M5 competition (with $s=1$) (Makridakis et al., 2022b).
-Lower values indicate better performance.
-To assess probabilistic accuracy, we evaluated the quantiles produced using a scaled version of the Multi-Quantile Loss.
-The Scaled Quantile Loss, also known as (Scaled) Pinball Loss, and the Scaled Multi-Quantile Loss are defined as:
-$\text{SQL}(\alpha)=\frac{\frac{1}{h}\sum_{t=n+1}^{n+h}\Big{(}\alpha\cdot(y_{t}-\hat{q}_{t}^{(\alpha)})\cdot\mathbb{I}_{\hat{q}_{t}^{(\alpha)}\leq y_{t}}+(1-\alpha)\cdot(\hat{q}_{t}^{(\alpha)}-y_{t})\cdot\mathbb{I}_{\hat{q}_{t}^{(\alpha)}>y_{t}}\Big{)}}{\frac{1}{n-s}\sum_{t=s+1}^{n}|y_{t}-y_{t-s}|}\mbox{,}$ (9)
-$\text{SMQL}=\frac{1}{|\mathcal{Q}|}\sum_{\alpha\in\mathcal{Q}}\text{SQL}(\alpha)\mbox{,}$ (10)
-These proper scoring rules assess forecast distribution accuracy (Kolassa, 2016).
-Moreover, the SMQL was the official metric of the M5 Uncertainty competition (weighted and with $s=1$) (Makridakis et al., 2022c).
-For this study, we adopted a cloud computing machine NC6s_v3 hosted on Microsoft Azure, with Linux Ubuntu 24 operating system, 1 Graphical Processing Unit, 6 Computing Processing Units, 112GB of memory.
 4 Results and discussion
 This section discusses the empirical findings of our study, highlighting the interplay between accuracy and stability across different retraining strategies.
-Table 2: Overall forecasting performance and stability for each method across datasets based on the baseline retraining scenario, $r=7$ for M5 and $r=1$ for VN1.
-Minimum values per column are highlighted in bold.
-Method M5 VN1
-RMSSE SMQL sMAPC SMQC RMSSE SMQL sMAPC SMQC
-LR 0.777 0.267 0.188 0.110 6.549 4.286 0.085 1.560
-RF 1.868 1.220 0.092 0.729
-XGBoost 0.755 0.245 0.075 0.100 1.890 1.503 0.146 0.629
-LGBM 0.771 0.246 0.054 0.118 3.542 2.318 0.121 1.484
-CatBoost 0.947 0.277 0.118 0.156 5.762 4.233 0.320 2.276
-MLP 0.821 0.263 0.500 0.083 1.543 0.806 0.458 0.219
-LSTM 1.913 0.892 0.558 0.242
-TCN 0.865 0.270 0.499 0.083 1.913 0.891 0.600 0.241
-NBEATSx 0.815 0.262 0.547 0.084 1.698 0.991 0.596 0.765
-NHITS 0.828 0.266 0.566 0.087 1.699 0.992 0.605 0.755
-Ens2A 0.757 0.243 0.059 0.108 1.472 0.826 0.503 0.441
-Ens3A 0.757 0.248 0.089 0.107 1.517 0.867 0.514 0.539
-Ens4A 0.758 0.240 0.095 0.099 1.524 0.926 0.167 0.545
-Ens5A 0.763 0.241 0.099 0.095 1.544 1.013 0.164 0.540
-Table 2 summarizes the performance and stability of all forecasting methods evaluated in this study, across both the M5 and VN1 datasets.
-For each method, the table reports four key metrics per dataset: RMSSE (Root Mean Squared Scaled Error) to evaluate point forecast accuracy, SMQL (Scaled Multi-Quantile Loss) to assess probabilistic forecast performance, sMAPC (symmetric Mean Absolute Percentage Change) to compare the stability in point forecast settings, and SMQC (Scaled Multi-Quantile Change) to understand the stability in probabilistic terms.
-Overall, the models evaluated achieved better stability results on the M5 dataset than on VN1.
-Several factors may account for this difference, such as the larger dataset size and a higher frequency of the M5 time series, the availability of rich external regressors (e.g., promotions, special events), and the existence of well-established benchmark hyperparameter settings, many of which were unavailable for VN1 at the time of the analysis.
-Moreover, in terms of point forecasting, machine learning models are more stable than deep learning methods on both datasets, implying that the former are less sensitive to frequent updates.
-On the contrary, in terms of probabilistic forecasting, deep learning architectures show less instability, meaning that these approaches can be a better option when retraining is performed frequently.
-In general, ensembling seems to be a good solution to produce stable predictions when compared to base models.
-However, larger ensembles do not necessarily imply more stability, and it is not clear whether model diversity may have an impact on ensembles stability of predictions.
-Figure 1 shows the point forecast stability of each model along the different retraining scenarios for the M5 and the VN1 datasets.
-To allow comparisons, we display the results in relative terms with respect to the baseline scenarios, that is $r=7$ for M5 and $r=1$ for VN1.
+Figure 1: RMSSE and sMAPC results for each global method and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M4 and M5 datasets, and $r=1$ for the VN1 dataset.
+Figure 1 shows the point forecast accuracy (on the left) and the point forecast stability (on the right) of each model  [^4] For the long computing time required, the Random Forest model has not been trained on the M5 dataset, while the LSTM model has not been adopted on both the M4 and M5 datasets.
+Moreover, the CatBoost method has not been used on the M4 dataset since too few categorical variables are available for this dataset.
+along the different retraining scenarios for the M4, M5, and VN1 datasets.
+To allow comparisons, we display the results in relative terms with respect to the baseline scenarios, that is $r=7$ for M4 and M5, and $r=1$ for VN1.
+In terms of forecast stability, the models’ behavior is similar across the different datasets.
 As expected, the sMAPC profiles are non-increasing functions of the retraining scenario.
-The stability remains practically the same across update frequencies or even improves as the retraining period increases.
-Indeed, moving from a high to a low retraining scenario, the stability of most global models improves compared to the baseline 444 Some models, like LR and MLP, show slightly decreasing stability (higher sMAPC) with an increasing retraining period.
-This can be due to many factors (e.g.
-sensitivity to small data changes, lack of robustness to stale parameters), but further investigation is necessary.
+The stability remains practically the same, or even improves, as the retraining period increases.
+Indeed, moving from a high to a low retraining scenario, the stability of most global models improves compared to the baseline.
+In particular, for some models (i.e.
+CatBoost, LSTM, and TCN), the stability improvement is strong and consistent, reaching very low levels.
+These results imply that less frequent retraining may have a huge impact on the point forecast stability of global models.
+This can be explained by the fact that, when a forecasting origin is updated, and the model is retrained on the new data, the model’s predictions may differ consistently (even using the cross-learning approach) from those produced before the update.
+Therefore, less retraining is often synonymous with more stability.
+The statistical tests (see the Supplementary material) confirm the above results, showing that most periodic retraining scenarios are statistically different to the continuous retraining strategy in terms of point forecast stability at the 5% level.
+Similar considerations arise for the point forecast accuracy: the RMSSE profiles of the global models remain remarkably stable across retraining scenarios, sometimes displaying even modest improvements in accuracy when retrained less frequently (as is the case of deep learning models for the M4 dataset).
+In particular, under low periodic retraining schemes, the performance of most global models is statistically indistinguishable from the baseline configuration, and even under extreme settings (such as the no-retraining scenario) any degradation in accuracy remains below 5%.
+Overall, these findings show that global models can preserve, and in some cases even enhance, their forecasting accuracy over extended retraining intervals.
+Hence, frequent retraining may introduce unnecessary variability without delivering tangible gains in both predictive performance and forecast stability.
+Figure 2: SMQL and SMQC results for each method and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M4 and M5 datasets, and $r=1$ for the VN1 dataset.
+In a similar fashion, Figure 2 summarizes the relative performance (on the left) and stability (on the right) from a probabilistic forecasting perspective.
+In this context, we observe that, for daily data, the stability paths are constant across almost every retraining scenario, meaning that less frequent updates neither harm nor improve the probabilistic forecasting stability of most global models  [^5] There are a few spikes in the SMQC profiles of some models.
+These can be due to different factors (e.g., inappropriate hyperparameter values).
+Nevertheless, the profiles are stable overall.
 .
-In particular, for some models (i.e.
-CatBoost, LSTM, and TCN), the instability reduction is strong and consistent, reaching very low levels.
-These results imply that less frequent retraining may have a huge impact on the point forecast stability of global models.
-This can be explained by the fact that, when a forecasting origin is updated and the model is retrained on the new data, the model’s predictions may differ consistently (even using the cross-learning approach) from those produced before the update.
-Therefore, less retraining is often synonymous with more stability.
-Figure 1: sMAPC results for each method and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
-Figures 2, and 3 statistically confirms the above results for the M5 and VN1 datasets.
-It is clear that most periodic retraining scenarios are statistically different to the continuous retraining strategy in terms of point forecast stability at the 5% level.
-Figure 2: M5 Friedman-Nemenyi test results based on sMAPC.
-Figure 3: VN1 Friedman-Nemenyi test results based on sMAPC.
-In a similar fashion, Figure 4 summarizes the relative stability in a probabilistic forecasting setting.
-In this context, we observe that, for the M5 dataset, the stability (as measured by Scaled Multi Quantile Change) paths are constant across almost every retraining scenario, meaning that less frequent updates neither harm nor improve the probabilistic forecasting stability of most global models 555 There are a few strange spikes in the SMQC profiles of some models.
-These can be due to different factors (e.g.
-inappropriate hyperparameter values), but further investigation is necessary.
+For weekly data, instead, we observe some differences among the models.
+Some of them show constant improvements in stability as the retraining scenario increases, while others do not exhibit particular enhancements or deteriorations in most periodic retraining intervals.
+When shifting the focus from point to probabilistic forecasting, the impact of retraining frequency on accuracy becomes more pronounced.
+Unlike point forecasts, probabilistic accuracy is more sensitive to longer retraining intervals.
+In several cases, reducing the retraining frequency leads to a gradual deterioration in probabilistic performance, indicating that the estimation of predictive uncertainty is more affected by outdated model parameters than point predictions.
+Nevertheless, these effects remain limited for low periodic retraining scenarios, where differences in probabilistic accuracy are typically negligible (below 2%).
+Importantly, this sensitivity is not universal: some models exhibit clear accuracy losses, while others maintain stable probabilistic performance or even improve.
+Overall, these findings suggest that accuracy uncertainty estimates benefit more from regular updates, yet do not strictly require very frequent retraining to remain reliable.
+Figure 3: RMSSE, sMAPC, SMQL, and SMQC results for ETS method on the M5 dataset for each retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$.
+To further contextualize the stability findings for global models, we also evaluated two classical local forecasting approaches: automatic ARIMA and automatic ETS  [^6] For the long computing time required, the ARIMA model has not been trained on the M5 dataset.
 .
-For the VN1 dataset, instead, we observe some differences among the models.
-Some of them show constant improvements in stability as the retraining scenario increases (i.e.
-NBEATS, NHITS, and CatBoost), while most of them do not exhibit any significant enhancements or deteriorations in the retraining intervals.
-For both datasets, these results are confirmed by the statistical tests on SMQC over the different retraining periods (see Supplementary materials).
-Figure 4: SMQC results for each method and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
-Overall, the results from Figures 1 and 4, supported by the Friedman-Nemenyi tests, indicate that reducing the retraining frequency of global models does not negatively impact (and may even enhance) forecast stability, for both point and probabilistic forecasts.
-When considered alongside the findings of Zanotti (2025b), this provides strong evidence against the practice of continuous retraining in global forecasting models.
-Lower retraining scenarios preserve both accuracy and stability, while offering substantial savings in computational resources.
-Figure 5 shows the evolution of point forecast stability, across different retraining scenarios for the various ensemble configurations.
-The results on the M5 dataset indicate that reducing retraining scenario may negatively affect forecast stability.
-However, this impact is below 1% for most retraining scenarios.
-Indeed, statistical tests (in Supplementary material) on the significance of these differences show that most retraining scenarios are not statistically different from each other, with periodic retraining strategies being at least as good as continuous retraining.
-The results on the VN1, instead, indicate that there exists a clear benefit in lowering the retraining scenario in terms of forecast stability.
-Notably, on the M5 dataset the smallest ensemble (Ens2A) consistently achieves the lowest sMAPC values, since it is obtained by combining XGBoost and LGBM, that where both the most accurate and the most stable global forecasting models on that set of data.
-On the contrary, larger ensembles (such as Ens4A and Ens5A) stability profiles tend to outperform smaller ones in the VN1 dataset, suggesting that when base models are less accurate, bigger ensembles may be necessary to improve forecast stability.
-Figure 6 further supports these observations in the probabilistic domain.
-First of all, probabilistic forecasts of ensembles are generally more stable than those of the base models across retraining frequencies, especially in M5, where SMQC variations are within 1 percentage point.
-In the VN1 dataset, the stability improvement is evident as the models are retrained less frequently.
-Then, from a probabilistic perspective, all the different ensembles perform similarly in terms of stability of the predictions, without showing any relevant benefit of having larger ensembles with respect to smaller ones.
-Most importantly, the effect of ensembling on the profiles of stability, as measured by sMAPC or SMQC, is that it greatly reduces the variations across the different retraining scenarios, reducing both positive and negative stability changes.
-Figure 5: sMAPC results for each ensemble and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
-Figure 6: SMQC results for each ensemble and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
-Figures 7 and 8 illustrate the fundamental trade-off between forecast accuracy and stability, comparing RMSSE with sMAPC and SMQL with SMQC, respectively.
-These plots reveal that high accuracy does not always guarantee high stability.
-In particular, several models achieve low RMSSE or SMQL scores but exhibit high instability, especially in VN1, highlighting their sensitivity to changes in the training data and their volatility across forecast origins.
-Nevertheless, it is the ensemble models that most consistently approach the optimality of stability and accuracy.
-In both the point and probabilistic domains, even very simple ensembles manage to control variance and mitigate instability without sacrificing much, if any, predictive performance.
-The overall picture that emerges from these figures reinforces the idea ensembling offers a practical and effective strategy to counteract the forecast instability.
-Figure 7: The accuracy-stability trade-off. RMSSE vs sMAPC for the baseline scenario, $r=7$ for the M5 dataset, and $r=1$ for the VN1 dataset.
-Figure 8: The accuracy-stability trade-off. SMQL vs SMQC for the baseline scenario, $r=7$ for the M5 dataset, and $r=1$ for the VN1 dataset.
-This set of results has clear practical implications: in operational forecasting systems, especially in domains such as retail and supply chain management, the use of ensembles combined with lower retraining scenarios may offer a robust forecasting process, balancing performance with stability.
+Both local models exhibit highly consistent behaviors across datasets and retraining scenarios (see Supplementary material).
+As an illustrative example, Figure 3 reports the results for ETS on the M5 dataset.
+The accuracy profiles for both point and probabilistic forecasting, measured by RMSSE and SMQL respectively, are strictly increasing functions of the retraining period, indicating a systematic deterioration in performance as retraining becomes less frequent.
+Conversely, as expected, the stability profiles, captured by sMAPC and SMQC, are strictly decreasing, showing that less frequent retraining leads to progressively more stable forecasts.
+This monotonic trade-off highlights an intrinsic property of local models: stability improvements are usually obtained at the cost of accuracy.
+In contrast, global models display markedly different dynamics, often maintaining both accuracy and stability over a wide range of retraining periods, and, in some cases, even improving along one dimension without compromising the other.
+This comparison suggests that global forecasting models are inherently more robust to reduced retraining, benefiting from cross-series learning that mitigates the sensitivity to individual time series updates, a property that local models, by construction, cannot exploit.
+Overall, the results from Figures 1 and 2, supported by the statistical tests, indicate that reducing the retraining frequency of global models does not negatively impact (and may even enhance) forecast stability, for both point and probabilistic forecasts.
+These results indicate that the accuracy–stability trade-off in global forecasting models is substantially milder than in local approaches.
+When considered alongside the findings of Zanotti (2025) on energy consumption, this provides strong evidence against the practice of frequent retraining in global forecasting models.
+Lower retraining scenarios preserve both accuracy and stability, while offering substantial savings in computational resources and energy consumption of forecasting systems.
+Figure 4: The optimal retraining frequency identified for each forecasting objective, and for each series in the datasets considering all the models.
+Finally, to jointly assess how accuracy- and stability-driven objectives translate into different retraining policies, Figure 4 provides a distributional view of the optimal retraining frequency across time series, in both point and probabilistic forecasting settings.
+Overall, the results indicate that continuous retraining is not required to achieve either accurate or stable forecasts when using global models.
+Across both point and probabilistic objectives, the retraining frequency can be safely reduced in favor of a periodic strategy, with stability-based criteria generally favoring slightly less frequent updates, in some cases even no retraining.
+However, compared to point forecasting, probabilistic objectives tend to favor slightly more frequent retraining, suggesting that accurate and stable uncertainty estimation requires a more regular model re-estimation.
+In practical terms, for daily data, a retraining interval of approximately 21–30 days emerges as a reasonable compromise across objectives, while for weekly data, retraining can be postponed usually up to 8 weeks without compromising performance.
+These retraining schedules strike a balance between the need to update models in response to potential distributional changes and the desire to limit unnecessary updates.
+Importantly, adopting such periodic retraining strategies can reduce the computational cost of forecasting by up to 75% Zanotti (2025), leading to more efficient and sustainable forecasting systems without sacrificing accuracy and also improving forecast stability.
 5 Conclusions
-This study provides a comprehensive empirical assessment of the forecast stability of global models, focusing on the effects of retraining across both point and probabilistic forecasting, in the context of retail demand.
-Our findings demonstrate that global forecasting models are able to produce very stable predictions, and that they can achieve even more stability under reduced retraining strategies, challenging the prevailing wisdom that frequent model updates are necessary to maintain forecast stability.
-In terms of point forecasting, the results consistently show that retraining less frequently does not harm and often improves forecast stability.
-This trend is evident in both the M5 and the VN1 datasets, where many models exhibit monotonic or flat stability profiles as the retraining interval increases.
-Probabilistic forecast stability, as measured by the newly defined Scaled Multi-Quantile Change (SMQC), shows a very similar behavior, indicating that the stability of probabilistic forecasts does not deteriorate as the models are retrained less frequently.
-These results are statistically validated via Friedman-Nemenyi tests and indicate that continuous retraining is often not the optimal strategy.
-Ensembling can be considered a good strategy to enhance forecast stability.
-Notably, ensemble models not only improve stability metrics but also smooth the variation in stability across retraining scenarios, further reinforcing their role as a stabilizing force in global forecasting systems.
-However, from our analysis, it is not clear whether the size and the diversity of the ensemble may have an impact on stability, and more research is needed in this direction.
-The joint analysis of retraining and ensembling reveals that the two approaches act as complementary levers for stabilizing forecasts.
-Less frequent retraining reduces abrupt changes in model behavior across forecasting origins, while ensembling mitigates the instability arising from any single model’s sensitivity to input changes.
-When combined, these strategies offer a robust forecasting setup that minimizes the instability of both point estimates and prediction intervals without sacrificing predictive accuracy.
-Figures comparing the accuracy–stability trade-off clearly show that ensembles lie closer to the optimality, offering an optimal balance between these competing objectives.
+This study investigates the role of retraining frequency in global forecasting models, with a particular focus on its implications for both forecast accuracy and forecast stability in point and probabilistic settings.
+Using three large-scale benchmark datasets (M4, M5, and VN1), we provided an extensive empirical analysis across multiple machine learning and deep learning architectures, complemented by a comparison with classical local forecasting methods.
+These results contribute to a more nuanced understanding of retraining as a design choice in large-scale forecasting systems, rather than as a default operational requirement.
+Across all datasets and modeling families, we find that reducing the retraining frequency does not harm and often improves forecast stability.
+This result holds for both point and probabilistic forecasts, as measured by the proposed Scaled Multi-Quantile Change (SMQC) metric.
+In particular, stability profiles are typically flat or monotonically improving as retraining becomes less frequent, indicating that frequent retraining often introduces unnecessary variability in forecasts without delivering tangible benefits.
+These results are statistically validated via statistical tests and indicate that continuous retraining is often not the optimal strategy.
+While point forecast accuracy is largely unaffected by reduced retraining, remaining stable or even improving in many cases, probabilistic accuracy shows a slightly higher sensitivity to retraining frequency.
+Less frequent updates may lead to modest degradations in distributional accuracy, although these effects are generally negligible for low periodic retraining scenarios and remain limited even under no-retraining regimes.
+Contrary to the common view of a strict trade-off between forecasting accuracy and stability, our results indicate that, in the context of global forecasting models, this tension is largely attenuated.
+Moreover, in many cases, reducing the retraining frequency does not require sacrificing predictive performance and may even contribute to both more stable and more accurate forecasts.
+This behavior contrasts with what is often observed in local modeling settings, where model re-estimation can lead to pronounced fluctuations in both accuracy and forecast revisions.
+For global models, the shared learning across series appears to act as a regularizing mechanism, making them less sensitive to frequent updates and allowing lower retraining frequencies to jointly improve stability while preserving, or even enhancing, also accuracy.
 From a practical standpoint, these findings have several implications.
-First, businesses that rely on large-scale forecasting systems, such as those in retail, inventory management, and supply chain planning can adopt lower retraining strategies not only to conserve computational resources, but also to increase operational stability, without compromising on forecast accuracy.
-Second, incorporating ensembles into global forecasting frameworks provides an effective hedge against model instability, thereby increasing the reliability of forecasts used in decision-making processes.
-Finally, these results support a shift in emphasis from pure accuracy to stability-aware model selection and evaluation, in those contexts that are more sensitive to forecast reviews.
+First, businesses that rely on large-scale forecasting systems, such as those in retail, energy, and supply chain planning can adopt lower retraining strategies not only to reduce computational resources and the associated costs of forecasting systems, but also to increase operational stability, without compromising on forecast accuracy.
+Indeed, our results strongly support low, periodic retraining strategies (i.e.
+every 30 days or every 8 weeks) in global forecasting systems.
+Finally, these results promote a shift in emphasis from pure accuracy to stability-aware model selection and evaluation, especially in those contexts that are more sensitive to forecast reviews.
 Nevertheless, some limitations of the present study should be acknowledged.
-While the analysis covers a broad range of global models and two state-of-the-art retail forecasting datasets, the generalizability of the findings to other domains or data types (e.g., macroeconomic indicators, financial series, or healthcare time series) remains an open question.
-Moreover, this work deliberately focuses on the stability of global models, leaving the analysis of stability in local models as an open avenue.
-A direct comparison between global and local approaches could therefore represent a valuable direction for future research.
+While the analysis covers a broad range of global models and three state-of-the-art forecasting datasets, exploring the effects of retraining newer methods and different datasets may be a concrete area for future research.
+Moreover, this work focuses on the stability of global models.
+A deeper analysis of local approaches could therefore represent a valuable direction for future research.
 Additionally, the study focuses solely on full retraining schemes and does not explore incremental or online learning strategies, which may offer a different trade-off between accuracy and stability.
-Future research could extend this work by incorporating such adaptive methods, assessing alternative ensembling techniques (e.g., weighted or stacking ensembles), and introducing business-oriented stability metrics that link forecast instability directly to operational cost or risk.
+Future research could extend this work by incorporating such adaptive methods and introducing business-oriented stability metrics that link forecast stability directly to operational cost or risk.
+Future research should also investigate the interplay between retraining frequency and hyperparameter optimization.
+Different models may exhibit heterogeneous responses to retraining strategies, depending on their hyperparameter configurations, and a deeper understanding of these interactions could enable the design of even more efficient and robust forecasting strategies.
 Furthermore, this study employed a single framework for generating probabilistic forecasts, namely Conformal Inference.
-While conformalized residuals provide a valid means for quantifying uncertainty, the resulting quantiles may primarily capture the stability of the conformal adjustment rather than that of the underlying predictive distributions.
-Future research could address this limitation by studying probabilistic stability in native probabilistic models, thereby enhancing the robustness and generalizability of the findings.
+Although conformalized residuals provide a valid means for quantifying uncertainty, future research could address this limitation by studying probabilistic stability in native probabilistic models, thereby enhancing the robustness and generalizability of the findings.
 Finally, further theoretical development of stability measures, particularly for probabilistic forecasts, may also contribute to a deeper understanding of how models behave across time and under varying data regimes.
-In conclusion, this work provides strong empirical evidence that frequent retraining is not necessary for maintaining stability in global forecasting models, and that ensembling is an effective strategy for achieving both stable and accurate forecasts.
-These findings contribute to a growing body of research promoting sustainable and robust forecasting practices in modern data-intensive environments.
+In summary, this work provides strong empirical evidence that frequent retraining is neither necessary nor desirable in many forecasting applications.
+By explicitly accounting for stability alongside accuracy, forecasting systems can be designed to be more robust and computationally sustainable.
 Supplementary material
-In this section, we provide tables and figures related to the empirical results of the M5 and VN1 datasets.
-Table 3 shows the composition of ensembles across the different datasets.
-The Tables 4, 5, 7, and 6 show the forecast instability of the different models and ensembles along the examined retrain scenarios for the M5 daily dataset.
-The Tables 8, 9, 11, and 10 show the forecast instability of the different models and ensembles along the examined retrain scenarios for the VN1 weekly dataset.
-Figures 9, and 10 show the results of the Friedman-Nemenyi test in the context of probabilistic instability for the M5 and VN1 datasets.
-Figures 11, 12, 13, and 14 show the results of the Friedman-Nemenyi test in the context of both point and probabilistic instability for the different ensemble models across retrain scenarios and for each dataset.
-Table 3: Composition of ensembles for the M5 and VN1 datasets.
-Dataset Ensemble Models Included
-M5 ENS2A XGBoost, LGBM
-ENS3A XGBoost, LGBM, LR
-ENS4A XGBoost, LGBM, LR, NBEATSx
-ENS5A XGBoost, LGBM, LR, NBEATSx, MLP
-VN1 ENS2A MLP, NBEATSx
-ENS3A MLP, NBEATSx, NHITS
-ENS4A MLP, NBEATSx, NHITS, RF
-ENS5A MLP, NBEATSx, NHITS, RF, XGBoost
+Model Learning rate Iterations Feature fraction
+Random Forest 0.03 1200 0.85
+XGBoost 0.03 1200 0.85
+LightGBM 0.03 1200 0.85
+CatBoost 0.03 1200
+Table 2: Main hyperparameter settings for machine learning forecasting models.
+Model Learning rate Epochs Scaler
+MLP 0.001 1000 Robust
+LSTM 0.001 1000 Robust
+TCN 0.001 1000 Robust
+NBEATS 0.001 1000 Robust
+NHITS 0.001 1000 Robust
+Table 3: Main hyperparameter settings for deep learning forecasting models.
 Method 7 14 21 30 60 90 120 150 180 364
-LR 0.188 0.188 0.188 0.189 0.189 0.190 0.190 0.192 0.193 0.196
-XGBoost 0.075 0.074 0.073 0.073 0.072 0.072 0.072 0.072 0.073 0.073
+ETS 1.606 1.760 1.925 2.035 2.510 3.041 3.341 3.704 4.113 5.449
+ARIMA 1.621 1.776 1.944 2.056 2.545 3.094 3.356 3.804 4.276 5.674
+LR 1.578 1.578 1.578 1.578 1.580 1.578 1.581 1.579 1.581 1.584
+RF 1.715 1.713 1.713 1.710 1.707 1.699 1.700 1.701 1.701 1.706
+XGBoost 18.125 18.115 18.140 18.142 18.144 18.116 18.130 18.026 18.088 18.112
+LGBM 15.467 15.452 15.450 15.435 15.413 15.393 15.386 15.282 15.261 15.034
+MLP 1.705 1.611 1.676 1.562 1.533 1.546 1.526 1.536 1.529 1.526
+TCN 2.280 2.043 1.997 1.815 1.727 1.733 1.699 1.758 1.761 1.730
+NBEATSx 1.625 1.521 1.494 1.475 1.468 1.471 1.469 1.473 1.475 1.498
+NHITS 1.629 1.548 1.522 1.506 1.493 1.494 1.496 1.496 1.499 1.503
+Table 4: M4 RMSSE values for each method and retrain scenario combination.
+Method 7 14 21 30 60 90 120 150 180 364
+ETS 0.014 0.010 0.008 0.007 0.005 0.004 0.003 0.003 0.002 0.000
+ARIMA 0.014 0.010 0.008 0.007 0.005 0.004 0.003 0.003 0.002 0.000
+LR 0.013 0.013 0.013 0.013 0.013 0.013 0.013 0.013 0.013 0.013
+RF 0.013 0.013 0.013 0.013 0.013 0.013 0.013 0.013 0.013 0.013
+XGBoost 0.052 0.052 0.052 0.052 0.052 0.052 0.052 0.052 0.052 0.052
+LGBM 0.123 0.123 0.123 0.123 0.123 0.122 0.122 0.122 0.122 0.121
+MLP 0.017 0.016 0.017 0.016 0.016 0.016 0.016 0.016 0.016 0.016
+TCN 0.021 0.017 0.016 0.015 0.015 0.015 0.015 0.015 0.014 0.015
+NBEATSx 0.015 0.015 0.015 0.015 0.015 0.015 0.015 0.015 0.015 0.016
+NHITS 0.016 0.015 0.015 0.015 0.015 0.015 0.015 0.015 0.015 0.015
+Table 5: M4 sMAPC values for each method and retrain scenario combination.
+Method 7 14 21 30 60 90 120 150 180 364
+ETS 0.419 0.467 0.513 0.549 0.684 0.864 0.948 1.078 1.230 1.563
+ARIMA 0.428 0.475 0.524 0.560 0.701 0.881 0.955 1.111 1.276 1.655
+LR 0.394 0.394 0.393 0.395 0.397 0.398 0.397 0.397 0.398 0.400
+RF 0.437 0.438 0.437 0.439 0.440 0.440 0.439 0.439 0.440 0.443
+XGBoost 2.279 2.288 2.293 2.309 2.337 2.365 2.394 2.405 2.447 2.636
+LGBM 2.455 2.473 2.480 2.511 2.564 2.624 2.680 2.707 2.793 3.096
+MLP 0.501 0.465 0.470 0.443 0.437 0.439 0.433 0.436 0.437 0.429
+TCN 0.652 0.583 0.577 0.531 0.508 0.532 0.506 0.528 0.521 0.507
+NBEATSx 0.466 0.426 0.415 0.410 0.409 0.409 0.410 0.407 0.409 0.421
+NHITS 0.466 0.435 0.423 0.419 0.414 0.414 0.414 0.412 0.414 0.415
+Table 6: M4 SMQL values for each method and retrain scenario combination.
+Method 7 14 21 30 60 90 120 150 180 364
+ETS 0.278 0.300 0.326 0.351 0.396 0.435 0.449 0.469 0.493 0.603
+ARIMA 0.279 0.301 0.328 0.347 0.397 0.439 0.451 0.482 0.496 0.593
+LR 0.262 0.262 0.262 0.261 0.263 0.264 0.266 0.266 0.268 0.276
+RF 0.351 0.351 0.351 0.352 0.352 0.353 0.354 0.353 0.353 0.357
+XGBoost 2.156 2.150 2.154 2.153 2.149 2.142 2.134 2.126 2.130 2.119
+LGBM 1.935 1.927 1.928 1.915 1.906 1.904 1.890 1.864 1.861 1.795
+MLP 0.371 0.359 0.369 0.346 0.349 0.353 0.354 0.357 0.350 0.360
+TCN 0.506 0.397 0.375 0.352 0.323 0.345 0.343 0.332 0.320 0.331
+NBEATSx 0.338 0.330 0.330 0.330 0.331 0.333 0.341 0.337 0.341 0.354
+NHITS 0.344 0.332 0.329 0.328 0.326 0.326 0.330 0.328 0.331 0.330
+Table 7: M4 SMQC values for each method and retrain scenario combination.
+Method 7 14 21 30 60 90 120 150 180 364
+ETS 0.747 0.751 0.756 0.759 0.769 0.776 0.778 0.774 0.784 0.782
+LR 0.759 0.759 0.759 0.759 0.758 0.758 0.758 0.758 0.759 0.758
+XGBoost 0.737 0.737 0.737 0.737 0.737 0.737 0.737 0.737 0.736 0.735
+LGBM 0.753 0.753 0.753 0.753 0.753 0.753 0.753 0.753 0.753 0.753
+CatBoost 0.917 0.929 0.924 0.928 0.921 0.936 0.925 0.930 0.934 0.926
+MLP 0.803 0.800 0.800 0.801 0.801 0.802 0.799 0.801 0.801 0.800
+TCN 0.844 0.844 0.844 0.844 0.844 0.844 0.844 0.843 0.843 0.841
+NBEATSx 0.797 0.796 0.796 0.807 0.797 0.797 0.797 0.797 0.797 0.796
+NHITS 0.811 0.808 0.817 0.808 0.808 0.806 0.807 0.807 0.807 0.805
+Table 8: M5 RMSSE values for each method and retrain scenario combination.
+Method 7 14 21 30 60 90 120 150 180 364
+ETS 0.104 0.070 0.054 0.045 0.026 0.018 0.012 0.013 0.007 0.000
+LR 0.188 0.188 0.188 0.188 0.189 0.190 0.190 0.191 0.192 0.196
+XGBoost 0.075 0.074 0.073 0.073 0.072 0.073 0.072 0.072 0.073 0.073
 LGBM 0.054 0.053 0.053 0.053 0.052 0.052 0.052 0.052 0.052 0.052
-CatBoost 0.118 0.043 0.059 0.037 0.029 0.020 0.018 0.018 0.018 0.019
-MLP 0.500 0.511 0.506 0.520 0.513 0.524 0.510 0.525 0.475 0.553
-TCN 0.499 0.209 0.174 0.086 0.055 0.043 0.031 0.017 0.017 0.003
-NBEATSx 0.547 0.541 0.527 0.561 0.528 0.536 0.523 0.536 0.515 0.412
-NHITS 0.566 0.548 0.558 0.534 0.565 0.537 0.533 0.520 0.512 0.564
+CatBoost 0.120 0.045 0.060 0.039 0.031 0.022 0.021 0.020 0.020 0.023
+MLP 0.499 0.511 0.505 0.519 0.512 0.524 0.510 0.525 0.475 0.552
+TCN 0.497 0.208 0.174 0.086 0.055 0.043 0.031 0.017 0.017 0.003
+NBEATSx 0.547 0.540 0.527 0.560 0.528 0.535 0.522 0.536 0.546 0.424
+NHITS 0.566 0.547 0.557 0.533 0.566 0.536 0.532 0.548 0.511 0.565
 Table 9: M5 sMAPC values for each method and retrain scenario combination.
 Method 7 14 21 30 60 90 120 150 180 364
-Ens2A 0.059 0.058 0.058 0.058 0.058 0.058 0.058 0.058 0.058 0.058
-Ens3A 0.089 0.089 0.089 0.089 0.089 0.089 0.090 0.090 0.091 0.092
-Ens4A 0.095 0.094 0.094 0.097 0.095 0.095 0.095 0.095 0.096 0.097
-Ens5A 0.099 0.098 0.098 0.101 0.099 0.099 0.099 0.099 0.100 0.101
-Table 5: M5 sMAPC values for each ensemble and retrain scenario combination.
+ETS 0.249 0.251 0.254 0.256 0.260 0.263 0.264 0.264 0.267 0.288
+LR 0.267 0.268 0.269 0.268 0.270 0.271 0.272 0.271 0.272 0.273
+XGBoost 0.245 0.246 0.247 0.247 0.249 0.250 0.252 0.251 0.251 0.253
+LGBM 0.246 0.247 0.247 0.247 0.248 0.249 0.250 0.249 0.250 0.251
+CatBoost 0.277 0.268 0.278 0.272 0.273 0.269 0.265 0.271 0.268 0.267
+MLP 0.263 0.264 0.265 0.266 0.269 0.271 0.271 0.272 0.275 0.276
+TCN 0.270 0.271 0.272 0.273 0.276 0.278 0.279 0.279 0.281 0.282
+NBEATSx 0.262 0.263 0.264 0.268 0.268 0.270 0.271 0.271 0.274 0.275
+NHITS 0.266 0.266 0.273 0.268 0.271 0.273 0.274 0.273 0.276 0.278
+Table 10: M5 SMQL values for each method and retrain scenario combination.
 Method 7 14 21 30 60 90 120 150 180 364
+ETS 0.105 0.105 0.105 0.104 0.103 0.102 0.102 0.101 0.103 0.102
 LR 0.110 0.110 0.110 0.111 0.111 0.111 0.112 0.112 0.112 0.113
@@
 Table 11: M5 SMQC values for each method and retrain scenario combination.
-Method 7 14 21 30 60 90 120 150 180 364
-Ens2A 0.108 0.108 0.108 0.108 0.108 0.108 0.108 0.108 0.108 0.107
-Ens3A 0.107 0.107 0.107 0.108 0.107 0.107 0.107 0.108 0.107 0.107
-Ens4A 0.099 0.099 0.099 0.101 0.100 0.100 0.099 0.100 0.099 0.099
-Ens5A 0.095 0.095 0.095 0.096 0.096 0.095 0.095 0.096 0.095 0.095
-Table 7: M5 SMQC values for each ensemble and retrain scenario combination.
 Method 1 2 3 4 6 8 10 13 26 52
-LR 0.085 0.084 0.085 0.082 0.083 0.086 0.086 0.085 0.084 0.082
-RF 0.092 0.090 0.089 0.089 0.088 0.088 0.089 0.089 0.089 0.090
-XGBoost 0.146 0.129 0.126 0.120 0.118 0.115 0.114 0.119 0.117 0.113
+ETS 1.886 1.910 2.033 1.967 2.032 2.050 2.146 2.398 2.113 2.036
+ARIMA 1.935 1.944 1.969 1.996 1.897 1.789 2.010 1.898 1.821 1.830
+LR 6.886 6.873 6.925 6.886 6.892 6.888 6.870 6.947 7.037 7.077
+RF 1.871 1.873 1.872 1.868 1.867 1.866 1.861 1.863 1.852 1.840
+XGBoost 1.898 1.871 1.872 1.865 1.842 1.851 1.809 1.833 1.851 1.797
+LGBM 3.540 3.520 3.471 3.282 3.414 3.732 3.415 3.337 3.357 3.543
+CatBoost 5.631 5.570 5.694 5.025 4.846 5.359 4.418 4.876 5.212 5.343
+MLP 1.554 1.483 1.477 1.472 1.472 1.475 1.480 1.480 1.467 1.455
+LSTM 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960
+TCN 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960
+NBEATSx 1.704 1.696 1.661 1.561 1.519 1.525 1.544 1.462 1.455 1.456
+NHITS 1.705 1.680 1.629 1.556 1.547 1.460 1.528 1.453 1.455 1.451
+Table 12: VN1 RMSSE values for each method and retrain scenario combination.
+Method 1 2 3 4 6 8 10 13 26 52
+ETS 0.156 0.114 0.093 0.070 0.054 0.039 0.032 0.034 0.014 0.000
+ARIMA 0.150 0.098 0.080 0.060 0.046 0.034 0.027 0.030 0.012 0.000
+LR 0.086 0.084 0.085 0.083 0.083 0.086 0.086 0.085 0.084 0.083
+RF 0.093 0.090 0.090 0.090 0.088 0.089 0.089 0.090 0.089 0.090
+XGBoost 0.146 0.129 0.125 0.120 0.118 0.115 0.114 0.119 0.116 0.113
 LGBM 0.121 0.110 0.111 0.106 0.104 0.105 0.105 0.104 0.104 0.105
-CatBoost 0.320 0.198 0.186 0.152 0.134 0.139 0.115 0.117 0.100 0.095
-MLP 0.458 0.469 0.468 0.475 0.443 0.483 0.438 0.490 0.469 0.452
-LSTM 0.558 0.301 0.203 0.126 0.101 0.044 0.055 0.048 0.032 0.013
-TCN 0.600 0.260 0.226 0.177 0.120 0.091 0.052 0.056 0.020 0.013
-NBEATSx 0.596 0.594 0.596 0.577 0.577 0.581 0.573 0.580 0.571 0.550
-NHITS 0.605 0.589 0.583 0.572 0.533 0.541 0.552 0.528 0.573 0.580
+CatBoost 0.312 0.197 0.185 0.152 0.134 0.139 0.115 0.117 0.101 0.095
+MLP 0.457 0.468 0.467 0.475 0.442 0.483 0.437 0.492 0.468 0.451
+LSTM 0.557 0.301 0.202 0.126 0.101 0.044 0.055 0.048 0.032 0.013
+TCN 0.599 0.259 0.226 0.177 0.120 0.091 0.052 0.056 0.020 0.013
+NBEATSx 0.594 0.592 0.594 0.576 0.576 0.581 0.572 0.580 0.571 0.549
+NHITS 0.605 0.587 0.581 0.571 0.531 0.540 0.550 0.527 0.574 0.581
 Table 13: VN1 sMAPC values for each method and retrain scenario combination.
-Method 7 14 21 30 60 90 120 150 180 364
-Ens2A 0.503 0.515 0.505 0.495 0.499 0.504 0.487 0.485 0.495 0.509
-Ens3A 0.514 0.526 0.526 0.509 0.509 0.514 0.510 0.512 0.511 0.512
-Ens4A 0.167 0.170 0.162 0.153 0.145 0.137 0.148 0.134 0.135 0.145
-Ens5A 0.164 0.162 0.154 0.145 0.138 0.131 0.138 0.129 0.129 0.135
-Table 9: VN1 sMAPC values for each ensemble and retrain scenario combination.
 Method 1 2 3 4 6 8 10 13 26 52
-LR 1.560 1.552 1.546 1.543 1.562 1.525 1.547 1.643 1.682 1.900
+ETS 1.003 1.032 1.077 1.084 1.115 1.133 1.159 1.249 1.196 1.259
+ARIMA 0.914 0.936 0.953 0.955 0.974 0.951 1.016 1.021 1.044 1.141
+LR 4.270 4.292 4.412 4.378 4.451 4.294 4.295 4.454 4.358 4.669
+RF 1.220 1.230 1.232 1.224 1.237 1.254 1.257 1.284 1.313 1.326
+XGBoost 1.503 1.496 1.493 1.505 1.483 1.488 1.466 1.451 1.515 1.479
+LGBM 2.318 2.292 2.230 2.157 2.230 2.317 2.262 2.362 2.220 2.268
+CatBoost 4.203 4.122 4.292 3.970 3.735 4.028 3.549 3.798 3.863 4.192
+MLP 0.806 0.800 0.796 0.795 0.803 0.805 0.807 0.819 0.824 0.847
+LSTM 0.892 0.895 0.899 0.903 0.910 0.917 0.924 0.935 0.957 1.003
+TCN 0.891 0.895 0.899 0.903 0.910 0.917 0.924 0.935 0.957 1.003
+NBEATSx 0.991 0.995 0.966 0.897 0.861 0.866 0.893 0.825 0.839 0.875
+NHITS 0.992 0.996 0.944 0.894 0.876 0.814 0.880 0.820 0.833 0.862
+Table 14: VN1 SMQL values for each method and retrain scenario combination.
+Method 1 2 3 4 6 8 10 13 26 52
+ETS 0.533 0.460 0.428 0.375 0.347 0.323 0.298 0.296 0.228 0.151
+ARIMA 0.463 0.396 0.345 0.304 0.291 0.260 0.255 0.235 0.182 0.132
+LR 1.559 1.551 1.545 1.536 1.549 1.525 1.544 1.632 1.674 1.894
 RF 0.729 0.732 0.735 0.737 0.735 0.738 0.743 0.734 0.731 0.734
 XGBoost 0.629 0.598 0.600 0.584 0.585 0.575 0.556 0.613 0.599 0.589
-LGBM 1.484 1.433 1.457 1.471 1.502 1.349 1.459 1.394 1.593 1.445
-CatBoost 2.276 1.861 1.769 1.582 1.547 1.579 1.333 1.581 1.302 1.243
+LGBM 1.480 1.416 1.449 1.471 1.502 1.349 1.455 1.394 1.593 1.445
+CatBoost 2.265 1.859 1.765 1.582 1.546 1.577 1.329 1.576 1.302 1.243
 MLP 0.219 0.281 0.278 0.278 0.282 0.275 0.266 0.285 0.267 0.283
@@
 Table 15: VN1 SMQC values for each method and retrain scenario combination.
-Method 7 14 21 30 60 90 120 150 180 364
-Ens2A 0.441 0.476 0.452 0.391 0.354 0.332 0.364 0.300 0.299 0.312
-Ens3A 0.539 0.562 0.507 0.430 0.389 0.324 0.398 0.312 0.309 0.322
-Ens4A 0.545 0.557 0.522 0.465 0.433 0.387 0.443 0.376 0.376 0.389
-Ens5A 0.540 0.545 0.519 0.472 0.448 0.410 0.453 0.407 0.405 0.416
-Table 11: VN1 SMQC values for each ensemble and retrain scenario combination.
+Figure 5: RMSSE and sMAPC results for each local method and retrain scenario combination in relative terms with respect to the baseline scenario, r = 7 for the M4 and M5 datasets, and r = 1 for the VN1 dataset.
+Figure 6: SMQL and SMQC results for each local method and retrain scenario combination in relative terms with respect to the baseline scenario, r = 7 for the M4 and M5 datasets, and r = 1 for the VN1 dataset.
+Figure 7: M4 Friedman-Nemenyi test results based on RMSSE.
+Figure 8: M4 Friedman-Nemenyi test results based on sMAPC.
+Figure 9: M4 Friedman-Nemenyi test results based on SMQL.
+Figure 10: M4 Friedman-Nemenyi test results based on SMQC.
+Figure 11: M5 Friedman-Nemenyi test results based on RMSSE.
+Figure 12: M5 Friedman-Nemenyi test results based on sMAPC.
+Figure 13: M5 Friedman-Nemenyi test results based on SMQL.
 Figure 14: M5 Friedman-Nemenyi test results based on SMQC.
+Figure 15: VN1 Friedman-Nemenyi test results based on RMSSE.
+Figure 16: VN1 Friedman-Nemenyi test results based on sMAPC.
+Figure 17: VN1 Friedman-Nemenyi test results based on SMQL.
 Figure 18: VN1 Friedman-Nemenyi test results based on SMQC.
-Figure 11: M5 Friedman-Nemenyi test results on ensembles based on sMAPC.
-Figure 12: M5 Friedman-Nemenyi test results on ensembles based on SMQC.
-Figure 13: VN1 Friedman-Nemenyi test results on ensembles based on sMAPC.
-Figure 14: VN1 Friedman-Nemenyi test results on ensembles based on SMQC.
+Figure 19: M4 Friedman-Nemenyi test results for local models.
+Figure 20: M5 Friedman-Nemenyi test results for local models.
+Figure 21: VN1 Friedman-Nemenyi test results for local models.